### v2

# Degradation-Aware All-in-One Image Restoration via Latent Prior Encoding

Abstract
Real-world images often suffer from spatially diverse degradations such as haze, rain, snow, and low-light, significantly impacting visual quality and downstream vision tasks. Existing all-in-one restoration (AIR) approaches either depend on external text prompts or embed hand-crafted architectural priors (e.g., frequency heuristics); both impose discrete, brittle assumptions that weaken generalization to unseen or mixed degradations. To address this limitation, we propose to reframe AIR as learned latent prior inference, where degradation-aware representations are automatically inferred from the input without explicit task cues. Based on latent priors, we formulate AIR as a structured reasoning paradigm: (1) which features to route (adaptive feature selection), (2) where to restore (spatial localization), and (3) what to restore (degradation semantics). We design a lightweight decoding module that efficiently leverages these latent encoded cues for spatially-adaptive restoration. Extensive experiments across six common degradation tasks, five compound settings, and previously unseen degradations demonstrate that our method outperforms state-of-the-art (SOTA) approaches, achieving an average PSNR improvement of 1.68 dB while being three times more efficient.
Refer to captionRefer to caption
Figure 1: Comparison of common all-in-one image restoration paradigms. Existing approaches depend on explicit task-specific guidance through manual prompts or predefined architectural biases. Our DAIR learns degradation-aware representations directly from degraded images through latent prior inference.
1 Introduction
Real-world images are frequently degraded by factors like haze, rain, snow, low-light, and motion blur Tian et al. (2025a); Jiang et al. (2025). These spatially varying degradations may occur alternately or even overlap, depending on environmental and capture conditions Guo et al. (2024). Such entangled degradation patterns significantly reduce visual quality, resulting in impaired downstream vision tasks Tian et al. (2025a); Jiang et al. (2025). Recent deep learning-based single-task methods Dong et al. (2020); Valanarasu et al. (2022); Chen et al. (2021); Cai et al. (2023); Wang et al. (2022); Zhang et al. (2017) have made significant progress in addressing individual restoration challenges. However, deploying separate task-specific (TS) networks for each degradation type is computationally expensive and impractical, driving interest in AIR frameworks Jiang et al. (2025).
Current AIR approaches can be broadly categorized into two paradigms. (1) Adaptive feature learning methods, such as ADAIR Cui et al. (2024) and AirNet Li et al. (2022), utilize hand-crafted frequency priors or architectural inductive biases to automatically differentiate degradation types. Although these methods circumvent manual specification, they inherently depend on pre-defined assumptions about degradation characteristics, which often fail to generalize to novel or compound corruptions Gao et al. (2024). (2) Prompt-based restoration methods, such as PromptIR Vaishnav et al. (2023) and DFPIR Tian et al. (2025a), rely on explicit, manually provided prompts to guide the restoration network by specifying degradation types. These methods offer flexible control but face a fundamental ”chicken-and-egg” dilemma: in real-world scenarios, the degradation type and location are rarely known beforehand, yet the network requires this information to perform effective restoration Jiang et al. (2025). Moreover, TS instructions constrain generalization, particularly in cases of mixed degradations or varying homogeneous degradations (e.g., different noise levels). As illustrated in Fig. 1, both existing AIR paradigms have limited robustness and practical applicability in unconstrained environments.
To overcome these limitations, we reframe AIR as a latent prior inference problem. Unlike existing methods Vaishnav et al. (2023); Cui et al. (2024); Tian et al. (2025a) that rely on explicit degradation prompting or predefined architectural inductive biases, we propose learning degradation-aware representations directly from the degraded image through a multi-level feature descriptor inspired by the variational autoencoder (VAE) Kingma & Welling (2013). Our learned prior eliminates the need for manual degradation hints and enables spatially adaptive restoration of diverse and unseen degradations. Guided by these learned priors, the proposed degradation-aware AIR (DAIR) framework incorporates a ”where, which, what” reasoning paradigm: it learns degradation-aware feature selection (which), localizes corrupted regions using spatial attention maps (where), and adaptively fuses multi-scale global representations (what). Our unified strategy substantially enhances the flexibility and generalizability of blind restoration methods.
Our contributions are:
  • We propose reframe AIR to directly learn latent degradation priors from the corrupted image, eliminating the need for external manual prompts (MP).
  • We propose a reasoning image restoration paradigm comprising: (1) latent priors for learning degradation-aware representations, enabling decisions on ”which” encoder features to utilize for reconstruction; (2) spatially-adaptive degradation map (DM) that integrate frequency-domain cues with efficient element-wise attention, providing interpretable and localized restoration guidance on ”where” to focus beyond implicit attention; (3) cross-modal fusion of structural and color cues with global degradation priors via adaptive scaling and shifting to determine ”what” content to reconstruct; and (4) a decoder with linear complexity performing explicit spatial reasoning for leverages 1-3 cues, we termed it 3WD (which–where–what decoding).
  • DAIR consistently outperforms SOTA on six common restoration tasks (e.g., snow, low-light), five compound degradations (e.g., haze + rain, low-light + haze + rain), and unseen degradations, achieving an average 1.68 dB PSNR gain. It also improves downstream tasks on images with unknown multi-type degradations, e.g., boosting YOLOv12-L Tian et al. (2025b) object detection by up to 7.40 mAP over SOTA AIR methods, highlighting strong generalizability (Fig. 1).
2 Related works
2.1 Image-to-image Translation
Early image restoration methods were designed to address individual TS settings, such as denoising Zhang et al. (2017); Pang et al. (2021); Zhang et al. (2018a), dehazing Dong et al. (2020); Qin et al. (2020), derainingJiang et al. (2020); Ren et al. (2019), low-light image enhancement (LLIE) Wei et al. (2018); Yi et al. (2023), or deblurring Cho et al. (2021); Nah et al. (2017). These task-dependent approaches achieved impressive results; their limited scope prevented generalization across diverse restoration challenges. Transformer-based models, such as Restormer Zamir et al. (2022) and Uformer Wang et al. (2022), have explored multiple degradation scenarios; however, they still require separate training for individual tasks. To address this, recent methods shift towards AIR, aiming to handle diverse degradations within a unified framework. AirNet Li et al. (2022) pioneered this field by introducing contrastive learning to extract degradation representations, guiding restoration for unknown corruptions. ADAIR Cui et al. (2024) recalibrate features using adaptive frequency statistics, enabling task-agnostic restoration. Recent methods, such as DiffUIR Zheng et al. (2024), have also explored generative techniques (i.e., latent diffusion Rombach et al. (2022)). These unified approaches mark a significant step forward, minimizing TS training and advancing toward generalizable AIR solutions.
2.2 Prompt-Guided All-in-one Restoration
Image-to-image translation-based method commonly fails in separating unique degradations Brooks et al. (2023); Jiang et al. (2025). To counter this, recent AIR Gao et al. (2024); Vaishnav et al. (2023); Tian et al. (2025a) methods have incorporated MP to provide prior knowledge of degradation cues. PromptIR Vaishnav et al. (2023) integrates text-based prompts into a transformer framework, leveraging degradation-specific semantic information to guide the restoration process. OneRestore Guo et al. (2024) introduced visual and text prompts within a transformer-based framework to encode TS information. Recently, DFPIR Tian et al. (2025a) introduced a degradation-aware feature perturbation method that utilizes CLIP-encoded Radford et al. (2021) text prompts to guide channel shuffling and attention masking, enabling unified restoration across diverse degradations. UHD-Processor Liu et al. (2025) introduced a VAE-based framework with progressive frequency learning and MP for ultra-high-definition image restoration. Notably, unlike UHD-Processor and other prompt-based methods, our method learns latent degradation representations directly from the input, enabling blind restoration without user guidance.
Figure 2: Overview of the proposed DAIR framework: (a) Overall architecture, (b) Latent prior encoding (which), (c) Degradation mapping (where), (d) Latent fusion (what), (e) 3WD restoration module. Best viewed with Zoom.
3 Method
Overall pipeline.
Fig. 2 illustrates the details of our framework. Given a degraded input image $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$ with unknown degradation $\mathcal{D}_{R}$, our framework learn to infer stage-wise latent codes $x\ell$ ($\ell=0,\dots,3$) and a bottleneck global descriptor $\mu$, serving as degradation priors. The encoder employs a two-branch architecture that separately processes multi-scale structural features (e.g., luminance $L_{\ell}$) and color features (e.g., chrominance $C_{\ell}$), enabling more effective representation of both spatial and chromatic information Sharif et al. (2025); Yan et al. (2025). At each stage, $L_{\ell}$ and $C_{\ell}$ interact with $x_{\ell}$ via Degradation-Aware Encoder Blocks (DAEB), leveraging degradation information into the feature extraction process. The branches are fused as $F_{\ell}=L_{\ell}+C_{\ell}$ and used as residual refined feature propagation to the decoder. Concurrently, a mapping block transforms $(L_{\ell},C_{\ell},x_{\ell})$ into a compact DM $g_{\ell}$, which is also forwarded to the decoder for explicit degradation localization guidance. At the bottleneck, a $\mu$-guided fusion block merges the deepest LC features with the global degradation latent using adaptive scaling and shifting Perez et al. (2018), resulting in $F_{\mu}=\gamma\odot[L_{3}+C_{3}]+\beta$. The decoder reconstructs the clean image from $F_{\mu}$, leveraging the degradation focus map $g_{\ell}$ and the selected features $F_{\ell}$ derived from latent priors. This architecture enables the model to automatically identify and restore a wide range of degradations without explicit user input or task specification.
3.1 Learning Degradation Reasoning with Latent Prior
We propose a multi-scale latent descriptor Kingma & Welling (2013) that directly infers continuous degradation representations from corrupted images without user-guided MP. Given $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$, the encoder extracts hierarchical feature maps $\{\mathbf{x}_{\ell}\}_{\ell=0}^{3}$ using convolutional layers and residual channel-attention Hu et al. (2018) blocks that emphasize degradation-relevant cues. At the deepest stage, multi-head self-attention Vaswani et al. (2017) refines the features to obtain the bottleneck representation $\mathbf{z}$. The VAE parameterizes a distribution with $\bm{\mu}=f_{\mu}(\mathbf{z})$ and $\log\bm{\sigma}^{2}=f_{\sigma}(\mathbf{z})$, sampling $\mathbf{z}_{reparam}=\bm{\mu}+\bm{\sigma}\odot\bm{\epsilon}$ where $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. The decoder reconstructs the input image from the sampled latent. This multi-objective training ensures that hierarchical features $\{\mathbf{x}_{\ell}\}$ capture degradation-specific representations at multiple scales, while bottleneck statistics $\bm{\mu}$ encode global degradation semantics. Both serve as degradation-aware latent cues to condition the restoration model:
$\text{VAE}(\mathbf{I}_{deg})=\{\mathbf{x}_{\ell}\}_{\ell=0}^{3},\;\bm{\mu}.$ (1)
3.1.1 Latent Prior Encoding
Unlike existing methods (PromptIR Vaishnav et al. (2023), DFPIR Tian et al. (2025a), CAPTNet Gao et al. (2024)) that inject prompts during decoding, we inject VAE-derived degradation priors during encoding. This enables our encoders to generate both degradation-aware and degradation-agnostic skip connections, providing richer guidance for reconstruction. We separately encode the luminance and chrominance (LC) components Smith (1978)($\mathbf{L}_{\text{deg}}$, $\mathbf{C}_{\text{deg}}$) of the degraded input using dedicated LC encoders Yan et al. (2025); Sharif et al. (2025) for better structural and chromatic information. Fig. 2 (b) illustrates our latent prior encoding. Two parallel encoders, $\mathcal{E}_{L}(\cdot)$ and $\mathcal{E}_{C}(\cdot)$, process these components across multiple scales using DAEB. At stage $\ell$, we injects latent degradation priors $\mathbf{x}_{\ell}$ via value modulation:
$Q_{\ell}$ $=\mathbf{f}_{\ell}W_{Q},\quad K_{\ell}=\mathbf{f}_{\ell}W_{K},\quad\tilde{V}_{\ell}=(\mathbf{f}_{\ell}W_{V})\odot\mathbf{x}_{\ell}$ (2)
where $\mathbf{f}_{\ell}\in\{\mathbf{f}^{L}_{\ell},\mathbf{f}^{C}_{\ell}\}$ are intermediate features. The attention mechanism produces degradation-aware representations as $\text{DAEB}(\mathbf{f}_{\ell},\mathbf{x}_{\ell})=\text{softmax}(Q_{\ell}K_{\ell}^{\top}/\sqrt{d})\tilde{V}_{\ell}$. Combining these components, the encoder outputs at each stage are:
$L_{\ell}=\mathcal{E}^{L}_{\ell}(\mathbf{L}_{\text{deg}},\mathbf{x}_{\ell}),\quad C_{\ell}=\mathcal{E}^{C}_{\ell}(\mathbf{C}_{\text{deg}},\mathbf{x}_{\ell})$ (3)
At each stage, latent-derived priors $\mathbf{x}_{\ell}$ modulate the value features after multi-head projection, enabling semantically refined propagation that preserves both degradation-specific and clean structural information. During decoding, these priors serve as selective feature utilization for effective reconstruction.
3.1.2 Learnable Degradation Map
Traditional skip connections propagate noise from degraded inputs to the decoder, without spatial awareness of the degradation Mao et al. (2016). We counter these limitations with learnable DM that encode spatially-varying degradation patterns for effective reconstruction (Fig. 2 (c)). At encoder stage $\ell$, our Degradation Mapping Block generates scale-specific maps $g_{\ell}$ from LC features $L_{\ell},C_{\ell}$ and latent prior encoding $\mathbf{x}_{\ell}$. We first fuse complementary LC information: $\mathbf{f}^{LC}_{\ell}=\text{Conv}_{3\times 3}(L_{\ell}+C_{\ell})$. To capture frequency-domain degradation characteristics, we enhance features with spectral information via Fast Fourier Transform Cooley & Tukey (1965): $\mathcal{F}(L_{\ell})=\text{FFT}(L_{\ell}-\mu(L_{\ell}))$, $\mathcal{F}(C_{\ell})=\text{FFT}(C_{\ell}-\mu(C_{\ell}))$. The magnitude and phase components are concatenated and projected:
$\mathbf{f}_{\ell}^{\text{freq}}=\text{ReLU}(\text{Conv}_{1\times 1}([|\mathcal{F}(L_{\ell})|\parallel|\mathcal{F}(C_{\ell})|\parallel\angle\mathcal{F}(L_{\ell})\parallel\angle\mathcal{F}(C_{\ell})])),\quad\tilde{\mathbf{f}}^{LC}_{\ell}=\mathbf{f}^{LC}_{\ell}+\mathbf{f}_{\ell}^{\text{freq}}$ (4)
For computational efficiency, we employ element-wise attention:
$Q_{\ell}$ $=\text{Conv}_{1\times 1}(\tilde{\mathbf{f}}^{LC}_{\ell}),\quad K_{\ell}=V_{\ell}=\text{Conv}_{1\times 1}(\mathbf{x}_{\ell})$ (5)
$\mathbf{A}_{\ell}$ $=\sigma(Q_{\ell}\odot K_{\ell}),\quad\mathbf{y}_{\ell}=\mathbf{A}_{\ell}\odot V_{\ell}$ (6)
where $\sigma$ is sigmoid activation and $\odot$ denotes Hadamard product. The DM is generated via residual refinement:
$g_{\ell}=\phi_{\text{map}}(\text{ReLU}(\text{Conv}_{3\times 3}(\mathbf{y}_{\ell}+\tilde{\mathbf{f}}^{LC}_{\ell})))$ (7)
where $\phi_{\text{map}}$ is a two-layer convolutional head producing $g_{\ell}\in\mathbb{R}^{1\times H_{\ell}\times W_{\ell}}$. This design captures spatial-frequency degradation patterns with linear complexity, while providing interpretable and localized restoration guidance.
3.1.3 Latent Encoded Fusion
At the encoder bottleneck, we leverage the deepest LC features $\mathbf{L}_{3},\mathbf{C}_{3}\in\mathbb{R}^{B\times 320\times\frac{H}{8}\times\frac{W}{8}}$, along with the latent global descriptor $\mu$. As shown in Fig. 2 (d), our fusion mechanism integrates three components: (i) intra-branch self-attention Sharif et al. (2025); Cai et al. (2023) for independent modality refinement, where $\hat{\mathbf{L}}_{3}=\text{MHSA}(\mathbf{L}_{3})$ and $\hat{\mathbf{C}}_{3}=\text{MHSA}(\mathbf{C}_{3})$; (ii) cross-branch attention Sharif et al. (2025) for complementary information exchange, where chrominance attends to luminance: $\mathbf{F}=\text{LayerNorm}\left(\text{MHCA}(\hat{\mathbf{C}}_{3},\hat{\mathbf{L}}_{3})+\hat{\mathbf{C}}_{3}+\hat{\mathbf{L}}_{3}\right)$, enabling the model to leverage structural and color information while preserving individual branch characteristics; and (iii) adaptive feature modulation conditioned on the global degradation prior $\mu$, combining LC features through adaptive scaling and shifting for TS feature modulation while maintaining underlying image characteristics. Unlike standard FiLM Perez et al. (2018) that applies uniform modulation, our approach makes degradation-aware reconstruction:
$[\bm{\gamma}_{\text{struct}},\bm{\beta}_{\text{color}}]$ $=\phi_{\mu}(\mu),\qquad\bm{\gamma}_{\text{struct}},\bm{\beta}_{\text{color}}\in\mathbb{R}^{B\times 320\times\frac{H}{8}\times\frac{W}{8}}$ (8)
$\mathbf{F}_{\mu}$ $=\mathbf{F}_{\text{LC}}\odot(1+\bm{\gamma}_{\text{struct}})+\bm{\beta}_{\text{color}}$ (9)
where $\phi_{\mu}$ serves as a content decision network (two $1\times 1$ convolutions with SiLU) that determines what structural and chromatic content should be reconstructed based on inferred degradation characteristics. Our identity-anchored scaling $(1+\bm{\gamma}_{\text{struct}})$ ensures content reconstruction decisions are made as learned adjustments around the original cross-modal features, completing our “what” reasoning component.
3.2 Degradation-Aware Reconstruction
The proposed decoder addresses our fundamental reasoning questions by reconstructing images through the 3WD module (Fig. 2 (e)). At each decoder stage $\ell$, given upsampled features $U_{\ell-1}$ and combined encoder features $E_{\ell}=L_{\ell}+C_{\ell}$, we compute linear attention projections:
$Q_{\ell}=W_{Q}^{(\ell)}U_{\ell-1},\quad K_{\ell}=W_{K}^{(\ell)}g_{\ell},\quad V_{\ell}=W_{V}^{(\ell)}E_{\ell}$
Degradation-guided attention operates as $A_{\ell}=\sigma(Q_{\ell}\odot K_{\ell})$, enabling $g_{\ell}$ to directly modulate spatial restoration. Features are updated through:
$D_{\ell}=\phi_{\ell}(A_{\ell}\odot V_{\ell}+U_{\ell-1})$ (10)
Notably, our proposed 3WD achieves linear computational complexity of $O(HW)$ per stage, in contrast to the quadratic complexity $O(H^{2}W^{2})$ of standard attention mechanisms Vaswani et al. (2017). This efficiency stems from utilizing element-wise multiplication ($\odot$) between tensors of shape $H\times W\times C$, requiring exactly $HW\times C$ operations. We perceived final reconstruction with learned features and global residual: $\hat{\mathbf{I}}=\tanh(W_{\text{rec}}D_{1})+\mathbf{I}_{\text{deg}}$.
4 Experiments
4.1 Setup
Dataset and methods.
We evaluated our method under two degradation scenarios: common (non-overlapping) degradations and compound (overlapping) settings. For common degradations, we tested on standard tasks/datasets, including dehazing (SOTS Li et al. (2018)), deraining (Rain100L Fu et al. (2017)), desnowing (CCD Li et al. (2020)), deblurring Nah et al. (2017), deblurring (GoPro Nah et al. (2017)), denoising (DIV2K Agustsson & Timofte (2017)) with random noise levels ($\sigma\in[0,50]$) to improve generalization, and LLIE using the LSD dataset Sharif et al. (2025), captured in uncontrolled environments. In compound degradation, we employed the CDD dataset Guo et al. (2024), combining five degradations (e.g., haze+rain, low-light+haze+snow). Subsets like haze+snow and low-light+rain were reserved for testing generalization comparison on unseen degradations. Ablation and analysis, we combined four common tasks (denoising, desnowing, deraining, and LLIE) to analyze consistency.
We compared DAIR against transformer-based baseline methods (Uformer Wang et al. (2022), Restormer Zamir et al. (2022)), SOTA AIR approaches (ADAIR Cui et al. (2024), AIRNet Li et al. (2022)), prompting-based methods (DFPIR Tian et al. (2025a), PromptIR Vaishnav et al. (2023)), and diffusion-based latent enhancement (DiffUIR Zheng et al. (2024)). Single-task benchmarks included deraining (HDCWNet Zhu et al. (2021), TransWeather Valanarasu et al. (2022)), Retinex-based LLIE (RetinexNet Wei et al. (2018), Diff-Retinex Yi et al. (2023)).
Implementation
We first pre-train our latent encoder for 200,000 steps, combining six single and five compound (known) degradations using a composite loss (self-reconstruction + KL Kingma & Welling (2014) + discriminative latent regularizer Guo et al. (2023)). The resulting encoder is frozen and reused for all single, compound, and unseen degradation experiments without any fine-tuning. We only tune the main network for TS settings using the combination of L1 and SSIM losses for fair comparison. This network has been trained for 100,000 to 500,000 steps, depending on task complexity. We trained both the latent encoder and the main restoration network using the Adam optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, and learning rate $1\text{e-}4$. All training was conducted on a single NVIDIA RTX 3060 GPU with a batch size of 4, using randomly cropped $256\times 256$ patches as input. All baseline methods are trained using their official implementations and suggested hyperparameters to ensure fairness.
4.2 Results on Common Degradation
4.2.1 Multi-task (6D) Restoration
Table 1: Performance comparison across different image restoration tasks. Best results in bold red, second best underlined, and increment over best performing method highlighted in blue.
Method MP Params (M) GFLOPs  [^1] Computed on $256\times 256$ images. Lowlight Dehazing Denoising Desnowing Deblurring Deraining Average
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
Uformer Wang et al. (2022) 20.63 43.86 12.10/0.4897 29.15/0.9727 27.55/0.8201 24.80/0.8832 23.67/0.8074 28.74/0.8823 24.34/0.8092
Restormer Zamir et al. (2022) 26.13 141.75 15.33/0.6180 28.51/0.9738 20.41/0.5357 22.98/0.8585 25.41/0.8286 25.99/0.8164 23.10/0.7718
AIRNet Li et al. (2022) 8.93 301.27 8.87/0.3655 25.92/0.9663 28.87/0.8789 19.08/0.7300 22.81/0.8067 24.92/0.8027 21.75/0.7583
PromptIR Vaishnav et al. (2023) 35.59 158.14 14.73/0.5374 30.41/0.9580 23.96/0.6378 27.08/0.9169 26.04/0.8366 30.01/0.8992 25.37/0.7977
DiffUIR Zheng et al. (2024) 36.26 4400.00 10.76/0.4101 30.05/0.9797 20.25/0.6859 20.39/0.8014 25.03/0.8474 27.46/0.8917 22.32/0.7694
ADAIR Cui et al. (2024) 28.78 147.18 12.76/0.4974 30.15/0.9783 28.65/0.8715 25.86/0.8812 25.14/0.8200 29.56/0.8901 25.35/0.8231
DFPIR Tian et al. (2025a) 31.07 151.07 14.84/0.5232 28.29/0.9709 27.78/0.8341 21.74/0.7935 26.10/0.8271 28.02/0.8372 24.46/0.7977
DAIR (Ours) 18.08 45.65 16.87/0.6661 34.08/0.9864 29.12/0.8971 31.50/0.9570 26.82/0.8793 30.51/0.9117 28.15/0.8829
Improvement -10.70 -96.10 +1.54/+0.0481 +3.67/+0.0067 +0.25/+0.0182 +4.42/+0.0401 +0.72/+0.0427 +0.50/+0.0125 +2.78/+0.0598
Figure 3: Visual comparison for six common degradation settings. The proposed method produces consistent and plausible images compared to the existing methods.
Table 1 and Figure 3 illustrate DAIR’s performance across six common degradations. Achieving the highest average PSNR (28.15) and SSIM (0.8829), DAIR delivers significant improvements, including +5.64 PSNR and +0.0758 SSIM in desnowing, and +1.54 PSNR and +0.0481 SSIM in lowlight enhancement. Despite requiring only 45G FLOPs, DAIR outperforms computationally intensive methods like ADAIR Cui et al. (2024)(147G FLOPs) while operating without MP, ensuring scalability and adaptability. By leveraging degradation-aware latent priors, DAIR generates robust, high-quality restorations across diverse degradation types, making it ideal for real-world applications like autonomous systems and healthcare imaging.
4.2.2 Single-task Evaluation: LLIE and Desnowing
Table 2: Performance comparison on desnowing and LLIE tasks. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
(a) Desnowing (b) Low-light enhancement
Method PSNR↑/SSIM↑ Method PSNR↑/SSIM↑
HDCWNet Zhu et al. (2021) 29.25/0.9171 RetinexNet Wei et al. (2018) 14.19/0.5183
TransWeather Valanarasu et al. (2022) 23.30/0.7631 Diff-Retinex Yi et al. (2023) 15.38/0.5038
Uformer Wang et al. (2022) 28.86/0.9324 Uformer Wang et al. (2022) 13.05/0.5716
Restormer Zamir et al. (2022) 23.21/0.8686 Restormer Zamir et al. (2022) 15.78/0.6184
ADAIR Cui et al. (2024) 28.31/0.9300 ADAIR Cui et al. (2024) 15.19/0.6094
DAIR (Ours) 32.75/0.9632 DAIR (Ours) 16.30/0.6634
Improvement +3.50/+0.0308 Improvement +0.52/+0.0450
We evaluated DAIR on single restoration tasks to demonstrate its adaptability. Table 2 shows DAIR achieves 32.75 PSNR / 0.9632 SSIM for desnowing, outperforming HDCWNet Zhu et al. (2021) by +3.50 PSNR / +0.0308 SSIM. For LLIE, DAIR achieves 16.30 PSNR / 0.6634 SSIM, improving upon Restormer Zamir et al. (2022) by +0.52 PSNR / +0.0450 SSIM. By leveraging degradation-aware latent priors, DAIR can also adapt to varying degradation types without external MP.
4.3 Results on Compound Degradation
Table 3: Performance comparison across five compound degradation. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
Method Haze + Rain Low + Haze Low + Haze + Rain Low + Haze + Snow Low + Snow Average
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
Uformer Wang et al. (2022) 22.49/0.8889 21.37/0.8252 20.61/0.7838 18.64/0.6802 20.31/0.7150 20.68/0.7786
Restormer Zamir et al. (2022) 23.57/0.8825 21.44/0.7753 20.53/0.7265 18.46/0.6494 19.97/0.6664 20.79/0.7400
AIRNet Li et al. (2022) 18.58/0.7977 11.73/0.6905 13.58/0.6320 12.87/0.5898 13.76/0.5759 14.10/0.6572
PromptIR Vaishnav et al. (2023) 21.78/0.8818 20.63/0.8231 19.83/0.7705 18.42/0.6801 20.74/0.7070 20.28/0.7725
DiffUIR Zheng et al. (2024) 18.42/0.8022 14.28/0.7150 15.79/0.6568 15.33/0.6238 12.60/0.6292 15.28/0.6854
ADAIR Cui et al. (2024) 23.51/0.8966 22.78/0.8420 21.06/0.7908 20.46/0.7229 21.35/0.7215 21.83/0.7948
DFPIR Tian et al. (2025a) 20.81/0.8308 19.01/0.7784 18.82/0.7139 15.58/0.6307 19.70/0.6402 18.78/0.7188
DAIR (Ours) 25.25/0.9259 23.15/0.8541 22.03/0.8200 20.81/0.7613 21.79/0.7772 22.61/0.8277
Improvement +1.68/+0.0293 +0.37/+0.0121 +0.97/+0.0292 +0.35/+0.0384 +0.44/+0.0557 +0.78/+0.0329
Figure 4: Visual comparison for five compound degradations. The proposed DAIR can handle compound degradation and produce visually pleasing images.
We evaluated DAIR on compound degradation scenarios. As shown in Table 3, DAIR achieves the highest average PSNR of 22.61 and SSIM of 0.8277, outperforming the next-best method (ADAIR Cui et al. (2024)) by +0.78 PSNR / +0.0329 SSIM. Notable improvements include ”Haze + Rain” (+1.68 PSNR / +0.0293 SSIM) and ”Low + Snow” (+0.44 PSNR / +0.0557 SSIM). DAIR effectively restores overlapping degradations, demonstrating robust performance across complex scenarios. Fig. 4 highlights DAIR’s superior visual quality, producing artifact-free, natural images.
4.4 Latent prior Guided Unseen Task Restoration
Table 4: Performance comparison on unknown compound degradations. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
Method Haze + Snow Low + Rain Average
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
ADAIR Cui et al. (2024) 16.07/0.8130 21.67/0.7877 18.87/0.8003
PromptIR Vaishnav et al. (2023) 17.84/0.8032 21.15/0.7752 19.50/0.7892
DFPIR Tian et al. (2025a) 14.50/0.7577 21.52/0.7021 18.01/0.7299
DAIR (Ours) 18.65/0.8531 23.29/0.8299 20.97/0.8415
Improvement +0.81/+0.0499 +1.62/+0.0422 +1.47/+0.0523
The primary motivation for leveraging latent priors is to address unknown degradations while maintaining a clear distinction from known degradations. Fig. 5 highlights the restoration of unknown compound degradations, comparing DAIR with SOTA methods. Our approach effectively handles these degradations, as demonstrated by t-SNE visualization in Fig. 5 (b). Table 4 summarizes the results, showing that our method achieves notable improvements over existing approaches, with average gains of +1.47 dB PSNR and +0.0412 SSIM for unseen degradations.
Refer to caption Refer to caption Refer to caption Refer to caption (b)
Figure 5: DAIR performance on unseen tasks: (a) Visual results for unseen compound degradation (low-light + rain); (b) t-SNE embeddings showing separation of unseen and known degradations.
4.5 Degradation Map
Figure 6: Visualization and restoration with the proposed DM.
Our proposed DM effectively learns spatially correlated regions, enabling the decoder to focus on ”where” to reconstruct the degraded image. As shown in Fig. 6, the comparison with traditional skip connections highlights that our learnable approach precisely identifies affected areas and significantly outperforms conventional methods.
4.6 Restoration with 3WD
We evaluate DAIR with 3WD against self-attention (SA) Vaswani et al. (2017) and cross-attention (CA) Vaswani et al. (2017) baselines. As shown in Table 5, DAIR with 3WD consistently outperforms both baselines, achieving average improvements of +7.72 dB PSNR and +0.2605 SSIM, with significant gains in desnowing (+11.89 dB) and denoising (+8.24 dB). Additionally, DAIR operates with linear complexity $O(d^{2}\times H\times W)$ compared to the baselines’ quadratic $O(d\times H^{2}\times W^{2})$, delivering 85-257× speedup and 76-89% memory reduction while answering our reasoning cues.
Table 5: Performance comparison of proposed 3WD restoration with existing attention mechanism. Best results in bold red, second best underlined. Improvement over the best performing method highlighted in blue.
Method Complexity Denoising Desnowing Deraining Lowlight Average
(Speedup) PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
DAIR + SA Vaswani et al. (2017) $O(H^{2}W^{2})$ 18.90/0.6213 19.70/0.6738 22.94/0.6541 15.16/0.4348 19.18/0.5960
DAIR + CA Vaswani et al. (2017) $O(H^{2}W^{2})$ 20.87/0.6290 16.07/0.6469 20.90/0.6448 13.00/0.3854 17.71/0.5765
DAIR + 3WD (Ours) $O(HW)$ 29.11/0.8971 31.59/0.9581 30.21/0.9070 16.68/0.6638 26.90/0.8565
Improvement Linear scaling +8.24/+0.2681 +11.89/+0.2843 +7.27/+0.2529 +1.52/+0.2290 +7.72/+0.2605
4.7 Ablation Study
We conducted an ablation study to evaluate DAIR’s key components: Latent Priors Encoding (LPE) (”which”), Degradation Map (DM) (”where”), Latent fusion (LF) (”what”), and Restoration Block (3WD). As shown in Table 6, the inclusion of the proposed component progressively enhances performance, with DAIR achieving the highest average PSNR of 26.90 and SSIM of 0.8565. Notably, removing Latent Priors results in significantly worse performance (-4.23 PSNR / -0.0876 SSIM), highlighting its critical role in capturing detailed degradation descriptions, subsequently guiding the model in effectively segregating different types of degradations.
Table 6: Ablation study on key components. Tick () indicates the component is used, cross () indicates not used. Best results in bold red.
Method LP LF DM 3WD Denoise Desnowing Derain Lowlight Average
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
Base Model 23.42/0.7237 20.76/0.7696 22.70/0.7698 13.99/0.5111 20.22/0.6936
Base + LP 24.50/0.7417 20.75/0.7783 22.46/0.7658 14.23/0.5240 20.49/0.7025
Base + LP + LF 25.83/0.7799 21.21/0.8014 22.23/0.7545 15.04/0.5474 21.08/0.7208
Base + LP + LF + DM 26.80/0.8158 21.97/0.8442 22.79/0.7614 15.52/0.5540 21.77/0.7439
DAIR w/o Latent Prior 26.07/0.8855 26.83/0.8306 22.76/0.7741 15.00/0.5853 22.67/0.7689
DAIR 29.11/0.8971 31.59/0.9581 30.21/0.9070 16.68/0.6638 26.90/0.8565
Improvement +5.69/+0.1734 +10.83/+0.1885 +7.51/+0.1372 +2.69/+0.1527 +6.68/+0.1629
5 Conclusion
We propose DAIR, a unified framework for AIR that tackles the challenges of unknown and compound degradations by learning degradation characteristics directly from the degraded image itself, eliminating the need for manual text or visual prompts. Guided by a reasoning paradigm based on ”which features to use, where to focus while restoring, and what to restore”, DAIR enables spatially adaptive restoration through a lightweight decoder that effectively integrates all prior information. Extensive experiments demonstrate that DAIR surpasses SOTA methods across six common and five compound degradation scenarios, while robustly handling unseen cases, highlighting its scalability and generalizability. Further details on implementation, analysis, unseen cases, downstream vision tasks, and additional results are provided in the appendix.
APPENDIX
Appendix A Network Details
A.1 Motivation
A major challenge in all-in-one image restoration is encoding degradation-specific information in a compact yet generalizable form while recovering clean images $\mathbf{I}$ from degraded observations $\mathbf{I}_{deg}=\mathcal{D}_{R}(\mathbf{I})+\epsilon$, where degradations $\mathcal{D}_{R}$ may arise from diverse sources. Current methods rely on explicit degradation cues, limiting their ability to handle unknown or compositional degradations.
From an information-theoretic perspective, effective restoration requires learning the conditional distribution $p(\mathbf{I}|\mathbf{I}_{deg})$ by disentangling degradation-specific information from content-preserving features. We model this through a variational framework where degradation characteristics are encoded in latent variable $\mathbf{z}\sim p(\mathbf{z}|\mathbf{I}_{deg})$. The VAE objective $\mathcal{L}=-\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{I}_{deg})}[\log p_{\theta}(\mathbf{I}|\mathbf{z},\mathbf{I}_{deg})]+\beta\cdot\text{KL}(q_{\phi}(\mathbf{z}|\mathbf{I}_{deg})\|p(\mathbf{z}))$ learns meaningful degradation representations while ensuring generalization through prior regularization. Degradations exhibit domain-specific characteristics: let $\mathbf{I}_{deg}=[L,C]$ where $L$ represents luminance and $C$ chrominance components. Structural degradations primarily affect $L$ while color distortions manifest in $C$. Factorizing the posterior as $q_{\phi}(\mathbf{z}|\mathbf{I}_{deg})=q_{\phi_{L}}(\mathbf{z}_{L}|L)\cdot q_{\phi_{C}}(\mathbf{z}_{C}|C)$ enables specialized degradation modeling, reducing cross-domain interference and improving restoration through $p_{\theta}(\mathbf{I}|\mathbf{z}_{L},\mathbf{z}_{C},\mathbf{I}_{deg})$. We propose a hybrid VAE with separate LC encoders that automatically infers continuous degradation latent without explicit MP.
A.2 VAE Design
Architecture.
To effectively integrate the VAE into our framework, we design a hybrid U-Net encoder–decoder with multi-head self-attention Tian et al. (2025a) at the bottleneck. In particular, both the hybrid VAE and the reconstruction network incorporate the same spatial feature dimensions, enabling seamless interaction between latent representations and restoration features. The encoder progressively downsamples RGB images through four stages with channel dimensions $[40,80,160,320]$ using 3×3 convolutions (stride=2 for downsampling, padding=1). Each stage contains a Residual Attention Block He et al. (2016) combining spatial convolutions with channel attention Hu et al. (2018) (reduction ratio 8):
$\text{ResAttn}(\mathbf{f})=\mathbf{f}+\text{Conv}_{3\times 3}(\text{ReLU}(\text{Conv}_{3\times 3}(\mathbf{f})))\odot\sigma(\text{Conv}_{1\times 1}(\text{GAP}(\mathbf{f})))$ (11)
The bottleneck applies 2 layers of Multi-Head Self-Attention (4 heads, $d_{k}=80$) to the deepest features $\mathbf{x}_{3}\in\mathbb{R}^{H/8\times W/8\times 320}$:
$\mathbf{Q},\mathbf{K},\mathbf{V}$ $=\mathbf{x}_{3}\mathbf{W}^{Q},\mathbf{x}_{3}\mathbf{W}^{K},\mathbf{x}_{3}\mathbf{W}^{V}$ (12)
$\text{MHSA}(\mathbf{x}_{3})$ $=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{80}}\right)\mathbf{V}$ (13)
Latent parameters are computed as $\bm{\mu},\log\bm{\sigma}^{2}=\text{Conv}_{1\times 1}(\text{MHSA}^{2}(\mathbf{x}_{3}))$ with latent dimension 320 and reparameterization $\mathbf{z}=\bm{\mu}+\bm{\sigma}\odot\bm{\epsilon}$. The decoder mirrors the encoder using transposed convolutions (stride=2) and identical ResAttn blocks, producing reconstructions $\hat{\mathbf{I}}_{deg}=\tanh(\text{Conv}(\text{Decoder}(\mathbf{z})))$.
Objective Function.
We train the complete VAE architecture using a multi-component loss function:
$\mathcal{L}_{\text{VAE}}=\mathcal{L}_{\text{recon}}+\beta\mathcal{L}_{\text{KL}}+\lambda_{\text{con}}\mathcal{L}_{\text{SupCon}}$ (14)
where $\beta$ and $\lambda_{\text{con}}$ are balancing hyperparameters for different loss components to stabilize VAE training. We set $\beta=0.3$ to provide strong KL regularization ensuring proper latent space structure, and $\lambda_{\text{con}}=0.01$ to enable weak supervision for degradation separation at the deepest encoder level (pre-latent space).
The reconstruction loss employs L1 distance to ensure pixel-level fidelity:
$\mathcal{L}_{\text{recon}}=\|\mathbf{I}_{deg}-\hat{\mathbf{I}}_{deg}\|_{1}$ (15)
where $\mathbf{I}_{deg}$ is the input degraded image and $\hat{\mathbf{I}}_{deg}=D_{\theta}(\mathbf{z})$ is the reconstructed output from the decoder.
The KL divergence Kullback & Leibler (1951) term regularizes the latent distribution:
$\mathcal{L}_{\text{KL}}=\text{KL}(q_{\phi}(\mathbf{z}|\mathbf{I}_{deg})\|\mathcal{N}(\mathbf{0},\mathbf{I}))$ (16)
This loss term encourages the learned latent distribution $q_{\phi}(\mathbf{z}|\mathbf{I}_{deg})$ to be close to a standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, ensuring that the latent space has good sampling properties and prevents overfitting.
To encourage discriminative latent representations, we incorporate weakly Supervised Contrastive Learning Guo et al. (2023) on the spatially-pooled encoder:
$\mathcal{L}_{\text{SupCon}}=-\sum_{i}\log\frac{\sum_{j\in P_{i}}\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{j}/\tau)}{\sum_{k\neq i}\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{k}/\tau)}$ (17)
where $\mathbf{z}_{i}$ are the normalized features, $P_{i}$ represents samples with the same degradation label as sample $i$, and $\tau=0.1$ is the temperature parameter. This loss pulls together samples from the same degradation category while pushing apart samples from different categories in the latent space, encouraging the encoder to learn discriminative representations for different types of image degradations.
A.3 DAIR Details
Architecture.
We employ a dual-branch encoder with separate LC processing. RGB inputs are decomposed as $L=0.299R+0.587G+0.114B$, C=RGBLC=RGB-L. Illumination and chrominance estimators utilize depthwise convolutions (kernel size=5, groups=4) to generate adaptive maps. Input embeddings are computed as $\mathbf{f}_{lum}=\text{LeakyReLU}(\text{Conv}_{3\times 3}(\mathbf{I}_{deg}\cdot\text{IlluMap}+\mathbf{I}_{deg}))$ and $\mathbf{f}_{chrom}=\text{LeakyReLU}(\text{Conv}_{3\times 3}(\mathbf{I}_{deg}\cdot\text{ChromMap}+\mathbf{I}_{deg}))$.
The encoder processes four levels with channel dimensions progressing as $40\to 80\to 160\to 320$ for both LC branches. Downsampling uses $\text{Conv}_{3\times 3}$ with stride=2 and padding=1, followed by LeakyReLU (slope=0.1). The decoder mirrors this structure using $\text{ConvTranspose}_{2\times 2}$ (stride=2) for upsampling $320\to 160\to 80\to 40$. Final reconstruction applies $\text{Conv}_{3\times 3}(\text{decoder\_out})+\mathbf{I}_{deg}$ with tanh activation. All convolutions use padding=1 for embedding layers and one attention block per level.
Objective Function.
The restoration network loss combines L1 reconstruction loss with SSIM-based perceptual loss:
$\mathcal{L}_{\text{recon}}=\|\hat{\mathbf{I}}-\mathbf{I}\|_{1}+\lambda_{\text{ssim}}\left(1-\text{SSIM}\left(\frac{\hat{\mathbf{I}}+1}{2},\frac{\mathbf{I}+1}{2}\right)\right)$ (18)
where images are normalized from [1,1][-1,1] to $[0,1]$ range for SSIM computation, and $\lambda_{\text{ssim}}=1.0$. The L1 loss ensures pixel-level fidelity while SSIM preserves perceptual quality and structural information. This combination effectively handles both fine-grained details and global image structure during restoration.
A.4 Training Details
Algorithm 1 details the complete training procedure. Notably, the VAE is trained only once using known degradation types, and the same pre-trained weights are utilized across all experiments, including single-task restoration, unseen degradation scenarios, and ablation studies, demonstrating the framework’s practical versatility in handling multi-degradation restoration tasks.
Algorithm 1 DAIR training for AIR
0: Degraded images $\{\mathbf{I}_{deg,i}\}$, clean images $\{\mathbf{I}_{i}\}$, labels $\{y_{i}\}$
0: Trained VAE $\theta_{\text{VAE}}$, restoration network $\theta_{\text{REST}}$
1: Initialize HybridVAE($\theta_{\text{VAE}}$), RestNet($\theta_{\text{REST}}$)
2: Set $T_{1}=200\text{K}$, $T_{2}=500\text{K}$, $\beta_{\max}=0.3$, $\lambda_{\text{con}}=0.01$
3:Phase 1: VAE Pretraining
4:for $t=1$ to $T_{1}$ do
5:  $\hat{\mathbf{I}}_{deg},\bm{\mu},\log\bm{\sigma}^{2},\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}\leftarrow\text{VAE}(\mathbf{I}_{deg})$
6:  $\hat{\mathbf{z}}_{3}\leftarrow\text{L2Norm}(\text{GAP}(\mathbf{x}_{3}))$ {Contrastive features}
7:  $\mathcal{L}_{\text{recon}}=\|\hat{\mathbf{I}}_{deg}-\mathbf{I}_{deg}\|_{1}$
8:  $\mathcal{L}_{\text{KL}}=-\frac{1}{2}\sum(1+\log\bm{\sigma}^{2}-\bm{\mu}^{2}-\bm{\sigma}^{2})$
9:  $\mathcal{L}_{\text{SupCon}}=\text{SupConLoss}(\hat{\mathbf{z}}_{3},y)$
10:  $\beta(t)=\min(\beta_{\max},\beta_{\max}\cdot t/T_{1})$ {KL annealing}
11:  $\mathcal{L}_{\text{VAE}}=\mathcal{L}_{\text{recon}}+\beta(t)\mathcal{L}_{\text{KL}}+\lambda_{\text{con}}\mathcal{L}_{\text{SupCon}}$
12:  Update $\theta_{\text{VAE}}$ via $\nabla_{\theta_{\text{VAE}}}\mathcal{L}_{\text{VAE}}$
13:end for
14:Phase 2: Restoration Network Training
15: Freeze($\theta_{\text{VAE}}$) {Fix VAE parameters}
16:for $t=T_{1}+1$ to $T_{1}+T_{2}$ do
17:  $\bm{\mu},\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}\leftarrow\text{VAE}(\mathbf{I}_{deg})$ {No gradients}
18:  $\hat{\mathbf{I}}\leftarrow\text{RestNet}(\mathbf{I}_{deg},\mathbf{x}_{0},\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\bm{\mu})$
19:  $\mathcal{L}_{\text{recon}}=\|\hat{\mathbf{I}}-\mathbf{I}\|_{1}+(1-\text{SSIM}(\hat{\mathbf{I}},\mathbf{I}))$
20:  Update $\theta_{\text{REST}}$ via $\nabla_{\theta_{\text{REST}}}\mathcal{L}_{\text{recon}}$
21:end for
22:return $\theta_{\text{VAE}},\theta_{\text{REST}}$
Appendix B Additional Experimental Results
B.1 Quantitative Comparison
B.1.1 Low-light
We utilize the LSD dataset Sharif et al. (2025), collected in uncontrolled low-light settings, offering diverse indoor and outdoor scenes under varying conditions. Table 7 compares image restoration methods across extreme lowlight (under 50 Lux) and lowlight scenarios (50-200 lux), divided into indoor and outdoor subsets, using PSNR and SSIM metrics. Our method, DAIR, consistently outperforms others, with the best results highlighted in bold red and the second-best underlined. DAIR achieves significant gains, improving up to +2.64 PSNR and +0.0737 SSIM in extreme lowlight, showcasing its robustness and superior generalization in challenging conditions.
Table 7: Performance comparison across lighting conditions. Best results in bold red, second best underlined.
Method Extreme Lowlight Lowlight
Indoor Outdoor Indoor Outdoor
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
Uformer Wang et al. (2022) 11.98/0.5278 13.24/0.4628 10.39/0.5535 12.79/0.4145
Restormer Zamir et al. (2022) 14.99/0.6343 16.35/0.5828 16.07/0.7121 13.90/0.5429
AIRNet Li et al. (2022) 8.51/0.3678 9.80/0.3248 8.13/0.4585 9.05/0.3107
PromptIR Vaishnav et al. (2023) 14.74/0.5902 16.10/0.5038 14.59/0.6404 13.48/0.4154
DiffUIR Zheng et al. (2024) 9.77/0.5185 11.86/0.3713 9.77/0.5185 11.29/0.3193
ADAIR Cui et al. (2024) 12.92/0.5456 14.46/0.4783 10.93/0.5576 12.72/0.4079
DFPIR Tian et al. (2025a) 13.32/0.5643 15.68/0.4815 15.63/0.6599 14.74/0.3871
DAIR (Ours) 17.63/0.7080 17.94/0.6486 17.13/0.7180 14.77/0.5898
Improvement +2.64/+0.0737 +1.59/+0.0658 +1.06/+0.0059 +0.87/+0.0469
B.1.2 Denoising
Our experiment assesses the generalization of image denoising models under variable noise conditions, addressing limitations of prior works that train on fixed noise levels, leading to overfitting. We retrained baseline methods and our DAIR model using random Gaussian noise on DIV2K Agustsson & Timofte (2017) and evaluated on unseen datasets, BSD100 Martin et al. (2001) and Urban100 Kim et al. (2016). Table 8 illustrate performance comparison across these datasets and noise levels ($\sigma$ = 15, 25, 50). Notably, baseline models showed inconsistent performance; Restormer struggled at higher noise, DiffUIR Zheng et al. (2024) excelled on BSD100 at $\sigma$=50 but failed elsewhere, on six-task common AIR settings. Notably, MP-based methods, such as PromptIR Vaishnav et al. (2023), underperformed in the absence of precise noise-level prompts. This inconsistent performance, particularly among MP-based methods, underscores their inability to effectively segregate degradations, even on homogeneous tasks. In contrast, DAIR consistently achieved higher scores, with improvements of up to +2.2 dB PSNR at high noise levels, highlighting its superior robustness and ability to effectively understand degradations, thereby validating the effectiveness of its latent prior encoding.
Table 8: Performance comparison across datasets and noise levels. Best results in bold red, second best underlined.
Method DIV2K BSD100 Urban100
$\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
Uformer Wang et al. (2022) 31.71/0.9265 29.16/0.8753 25.46/0.7351 33.28/0.9219 30.61/0.8635 33.28/0.9219 31.52/0.9387 29.13/0.8944 25.15/0.7664
Restormer Zamir et al. (2022) 27.38/0.8567 23.44/0.6825 16.62/0.3646 28.32/0.8385 23.73/0.6373 31.59/0.8937 28.11/0.8873 23.67/0.7304 16.52/0.4315
AIRNet Li et al. (2022) 32.91/0.9377 30.52/0.9095 27.07/0.8307 34.20/0.9333 31.83/0.8986 27.03/0.7006 33.19/0.9506 31.00/0.9324 27.46/0.8665
PromptIR Vaishnav et al. (2023) 28.84/0.8464 25.90/0.7306 21.45/0.5164 30.04/0.8295 27.03/0.7007 30.61/0.8636 29.23/0.8755 26.05/0.7734 21.23/0.5685
DiffUIR Zheng et al. (2024) 19.92/0.8065 15.95/0.6774 14.34/0.5435 21.78/0.7909 17.41/0.6917 31.99/0.9122 18.28/0.7967 14.86/0.6701 13.67/0.5496
ADAIR Cui et al. (2024) 32.51/0.9412 30.10/0.9075 26.80/0.8174 34.06/0.9354 31.59/0.8937 14.07/0.4531 32.74/0.9514 30.33/0.9233 26.72/0.8420
DFPIR Tian et al. (2025a) 22.13/0.8095 21.57/0.7585 20.43/0.6234 21.77/0.8164 21.47/0.7535 21.48/0.7533 22.36/0.8182 21.80/0.7759 21.48/0.7533
DAIR (Ours) 33.08/0.9541 30.62/0.9252 27.25/0.8555 34.48/0.9468 31.99/0.9122 31.83/0.8988 33.23/0.9623 30.82/0.9395 27.30/0.8796
Improvement +0.17/+0.0164 +0.10/+0.0157 +0.18/+0.0248 +0.28/+0.0135 +0.16/+0.0136 -0.16/-0.0134 +0.04/+0.0117 -0.18/+0.0071 -0.16/+0.0131
B.2 More Visual Results
Fig. 7 presents visual results for denoising, deraining, and dehazing tasks, comparing our method with ADAIR Cui et al. (2024), DFPIR Tian et al. (2025a), degraded inputs, and reference images. These examples demonstrate that our approach consistently achieves clearer and more accurate restorations, closely matching the reference images and outperforming baselines across all degradation types.
Figure 7: Visual comparisons of common image restoration, including denoising, deraining, and dehazing. Our method consistently delivers clearer and more accurate outputs, closely matching the reference images and outperforming ADAIR and DFPIR.
Fig. 8 presents more visual results for compound degradations (e.g., low-light + haze + rain, haze + rain), comparing our method with ADAIR Cui et al. (2024), DFPIR Tian et al. (2025a), degraded inputs, and reference images. Our approach consistently delivers clearer and more accurate restorations, outperforming baselines in challenging compound degradation settings.
Figure 8: Visual comparisons of compound image restoration results for challenging scenarios, including haze + rain, haze + snow, and low-light + haze + rain. Our method consistently delivers clearer and more accurate outputs, closely matching the reference images and outperforming ADAIR and DFPIR.
B.3 Inference Benchmarking
Table 9: Inference speed comparison across different image restoration models. Best results in bold red, second best underlined.
Model 128×128 256×256 512×512
Time (ms)↓ FPS↑ Throughput (MP/s)↑ Time (ms)↓ FPS↑ Throughput (MP/s)↑ Time (ms)↓ FPS↑ Throughput (MP/s)↑
Uformer Wang et al. (2022) 31.07 32.19 0.53 58.08 17.22 1.13 203.83 4.91 1.29
PromptIR Vaishnav et al. (2023) 65.62 15.24 0.25 180.31 5.55 0.36 707.17 1.41 0.37
ADAIR Cui et al. (2024) 71.01 14.08 0.23 191.50 5.22 0.34 732.94 1.36 0.36
DFPIR (with CLIP) Tian et al. (2025a) 70.08 14.27 0.23 184.33 5.43 0.36 714.20 1.40 0.37
DFPIR (dummy text) Tian et al. (2025a) 67.88 14.73 0.24 184.12 5.43 0.36 714.38 1.40 0.37
DAIR (Ours) 20.55 48.67 0.80 55.20 18.12 1.19 220.92 4.53 1.19
Table 9 presents the inference speed of image restoration models, evaluated on lower-mid hardware (RTX 3060) using full-precision weights (FP32). Our DAIR model consistently achieves the fastest processing times, highest FPS, and superior throughput across various image sizes, outperforming all baseline AIR models. Notably, DAIR matches or surpasses the single-task model like Uformer (43.86G) Wang et al. (2022), which is significantly lighter compared to the existing AIR methods. These results highlight DAIR’s efficiency, demonstrating its ability to deliver SOTA restoration quality while maintaining faster speed and practicality for real-world applications, even on modest hardware setups.
B.4 Downstream Vision Tasks (Object Detection)
A primary objective of AIR is enhancing downstream vision task performance Cui et al. (2024); Jiang et al. (2025), yet existing AIR methods lack benchmarking on such tasks due to insufficient annotated datasets. To address this critical gap, we created a comprehensive evaluation framework by annotating 3,000 test images across six common degradations (lowlight Sharif et al. (2025), deraining Fu et al. (2017), dehazing Li et al. (2018), desnowing Li et al. (2020), denoising Agustsson & Timofte (2017), and deblurring Nah et al. (2017)). Our annotation process leveraged the paired nature of these datasets by first annotating ground truth reference images and transferring these annotations to their corresponding corrupted counterparts, ensuring consistency and reliability. We then enhanced these degraded images using several baseline AIR restoration methods and systematically evaluated their performance using the YOLOv12-Large Tian et al. (2025b) object detection model. This novel benchmarking approach provides the first quantitative assessment of how different restoration techniques impact object detection performance, offering valuable insights for developing AIR methods optimized for real-world computer vision applications.
Table 10: Performance comparison of object detection models. Best results in bold red, second best underlined.
Method Average Precision (AP) AP by Object Size
IoU=0.50:0.95 IoU=0.50 IoU=0.75 Small Medium Large
Uformer Wang et al. (2022) 28.0 30.2 29.4 36.5 26.9 27.3
Restformer Zamir et al. (2022) 31.7 34.1 33.2 37.5 28.7 32.4
AIRNet Li et al. (2022) 30.2 32.7 31.8 35.8 26.4 30.3
PromptIR Vaishnav et al. (2023) 30.9 33.3 31.9 41.8 28.6 30.6
ADAIR Cui et al. (2024) 28.9 31.2 30.2 38.4 28.8 28.1
DFPIR Tian et al. (2025a) 27.9 30.0 29.3 26.6 25.5 29.8
DAIR (Our) 34.9 37.3 36.4 44.9 30.8 35.1
Improvement +3.2 +3.2 +3.2 +3.1 +2.0 +2.7
Table 10 presents a comprehensive performance comparison of various image restoration methods for object detection tasks. Our proposed DAIR method demonstrates superior performance across all evaluation metrics, achieving the highest Average Precision (AP) scores of 34.9% at IoU=0.50:0.95, 37.3% at IoU=0.50, and 36.4% at IoU=0.75, surpassing the second-best method (Restformer Zamir et al. (2022)) by a significant margin of 3.2 percentage points.DAIR also excels in detecting objects of varying sizes, achieving notably high AP scores on small (44.9%, outperforming PromptIR’s 41.8%), medium (30.8%), and large (35.1%) objects. The consistent improvements across all metrics highlight DAIR’s effectiveness in enhancing image quality for downstream object detection, with an average improvement of 3.2 percentage points over SOTA methods. These results demonstrate DAIR’s robust capability to address diverse degradation scenarios while preserving essential visual information for accurate object detection.
Appendix C Unseen Task Generalization
A key motivation behind the proposed latent-prior DAIR framework is its ability to generalize to diverse, previously unseen degradations. To validate this capability, we extensively evaluate DAIR under cross-domain and out-of-distribution scenarios, including underwater image enhancement Peng & Cosman (2019), medical image perceptual enhancement Uhlen et al. (2010), medical image denoising Rezvantalab et al. (2018), real-world denoising Abdelhamed et al. (2018), and unseen real-world low-light enhancement Sharif et al. (2025). Notably, all evaluations are performed using a DAIR model trained only on common degradations, without any TS fine-tuning or external guidance. In addition to the visual results, we summarized the quantitative score with no-reference metrics: NIQE Mittal et al. (2013), MUSIQ Ke et al. (2021), and LPIPS Zhang et al. (2018b).
C.1 Real-World Denoising
Table. 8 and Figure 9 illustrate the performance of the proposed method on real-world noisy images Abdelhamed et al. (2018). In well-exposed scenes, DAIR performs only denoising, preserving natural brightness and contrast. However, when inputs exhibit low-light characteristics (as commonly observed in SSID-like datasets Abdelhamed et al. (2018)) along with noise, the latent descriptor guides DAIR to jointly suppress noise and enhance brightness. Thus, DAIR adaptively performs denoising and enhancement without prompts or handcrafted rules.
Table 11: Image Quality Metrics for Real-world Noise Dataset. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 5.57 61.31 0.81
Enhanced (Our) 4.24 69.37 0.80
Improvement +1.33 +8.06 +0.01
Figure 9: Real-world denoising results. Left: low-light noisy scenes where DAIR jointly performs denoising and LLIE automatically, without user input. Right: well-lit, noisy scenes where DAIR performs only denoising.
C.2 Underwater Image Enhancement
We further evaluate DAIR on underwater images, a domain unseen during training. As shown in Figure 10, the model adapts to varying underwater degradations, including haze-like scattering (left examples) and low exposure (right examples). Despite never being trained on underwater data, DAIR enhances visibility and color fidelity, demonstrating strong cross-domain generalization driven by its latent prior representation. Table 12 further details DAIR’s performance on underwater image enhancement.
Figure 10: Underwater image enhancement. DAIR generalizes to haze-like and low-exposed underwater degradations, improving contrast and color balance without domain-specific training.
Table 12: Image Quality Metrics for Underwater Image Enhancement. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 3.83 53.35 0.70
Enhanced (Our) 3.78 54.05 0.70
Improvement +0.05 +0.70 +0.00
C.3 Unseen Low-Light Image Enhancement
We evaluate DAIR on real-world low-light images captured with unknown hardware and distributions (e.g., LSD-U Sharif et al. (2025)) unseen during training. Figure 11 and Table 13 demonstrate that the latent prior effectively encodes degradation properties, enabling robust out-of-distribution LLIE.
Figure 11: Unseen real-world low-light enhancement. DAIR can enhance illumination and structure in previously unseen low-light scenes without artifacts or over-enhancing the scenes.
Table 13: Image Quality Metrics for Real-world Lowlight Enhancement. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 4.47 62.33 0.81
Enhanced (Our) 4.14 62.51 0.81
Improvement +0.33 +0.18 +0.00
C.4 Medical Image Perceptual Enhancement
We also study cross-domain perceptual enhancement on microscopic medical images Uhlen et al. (2010), which often suffer from poor contrast and weak structural saliency Sharif et al. (2022). Despite having no exposure to this modality during training, DAIR enhances perceptual quality by improving local contrast and clarity while preserving natural textures (Figure 12). These results demonstrate that the latent prior generalizes to structurally distinct domains. Table 14 presents the corresponding quantitative improvements.
Figure 12: Perceptual enhancement of microscopic medical images. DAIR improves contrast and structural visibility in unseen medical imagery without TS training.
Table 14: Image Quality Metrics for Perceptual Enhancement. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 13.94 37.06 0.63
Enhanced (Our) 9.68 41.50 0.79
Improvement -4.26 +4.44 -0.16
C.5 Medical Image Denoising
Finally, we evaluate DAIR on medical image denoising using a dataset not seen during training, specifically dermatoscopic imagery Rezvantalab et al. (2018). As shown in Figure 13, DAIR effectively suppresses noise while preserving diagnostically relevant textures and edges, without requiring manual prompting. Table 15 illustrates the performance of DAIR on unseen medical image denoising.
Figure 13: Medical image denoising. DAIR reduces noise and preserves fine structural details in previously unseen medical data, operating without guidance or fine-tuning.
Table 15: Image Quality Metrics for Medical Noise Reduction. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 16.99 21.93 0.82
Enhanced (Our) 12.64 25.17 0.70
Improvement +4.35 +3.24 +0.12
Appendix D Limitation and Future Scope
Despite the significant improvement of our proposed DAIR method over existing approaches on both common and compound degradations, as well as its performance on unseen tasks, we observed certain limitations. Figure 14 illustrates some failure cases of our method on unseen tasks. Specifically, in extreme low-light conditions (under 5 lux), artifacts were observed in homogeneous spatial regions, such as cloud-free skies. Additionally, in underwater image enhancement, the method occasionally produced over-enhanced results in tricky scenes. These observations highlight areas that require further investigation, which we plan to address in future studies.
Moreover, the proposed method was tested on lower-mid desktop environments. In contrast, AIR demonstrates significant potential for deployment on edge hardware. As part of future work, we plan to optimize DAIR for edge hardware by employing techniques such as low-bit quantization and mixed-precision inference. Furthermore, integrating vision tasks with AIR for practical applications presents another exciting direction for future research on DAIR.
Figure 14: Example of failure cases. Left: visual artifacts on the spatial region in extreme low-light conditions, right: over-enhancement of the underwater image.

### v3

@@
 Extensive experiments across six common degradation tasks, five compound settings, and previously unseen degradations demonstrate that our method outperforms state-of-the-art (SOTA) approaches, achieving an average PSNR improvement of 1.68 dB while being three times more efficient.
+Code will be released upon publication.
 Refer to captionRefer to caption
@@
 1 Introduction
-Real-world images are frequently degraded by factors like haze, rain, snow, low-light, and motion blur Tian et al.
-(2025a); Jiang et al.
-(2025).
-These spatially varying degradations may occur alternately or even overlap, depending on environmental and capture conditions Guo et al.
-(2024).
-Such entangled degradation patterns significantly reduce visual quality, resulting in impaired downstream vision tasks Tian et al.
-(2025a); Jiang et al.
-(2025).
-Recent deep learning-based single-task methods Dong et al.
-(2020); Valanarasu et al.
-(2022); Chen et al.
-(2021); Cai et al.
-(2023); Wang et al.
-(2022); Zhang et al.
-(2017) have made significant progress in addressing individual restoration challenges.
-However, deploying separate task-specific (TS) networks for each degradation type is computationally expensive and impractical, driving interest in AIR frameworks Jiang et al.
-(2025).
+Real-world images are frequently degraded by factors like haze, rain, snow, low-light, and motion blur (tian2025degradation; jiang2025survey).
+These spatially varying degradations may occur alternately or even overlap, depending on environmental and capture conditions (guo2024onerestore).
+Such entangled degradation patterns significantly reduce visual quality, resulting in impaired downstream vision tasks (tian2025degradation; jiang2025survey).
+Recent deep learning-based single-task methods (dong2020multi; valanarasu2022transweather; chen2021all; cai2023retinexformer; wang2022uformer; zhang2017beyond) have made significant progress in addressing individual restoration challenges.
+However, deploying separate task-specific (TS) networks for each degradation type is computationally expensive and impractical, driving interest in AIR frameworks (jiang2025survey).
 Current AIR approaches can be broadly categorized into two paradigms.
-(1) Adaptive feature learning methods, such as ADAIR Cui et al.
-(2024) and AirNet Li et al.
-(2022), utilize hand-crafted frequency priors or architectural inductive biases to automatically differentiate degradation types.
-Although these methods circumvent manual specification, they inherently depend on pre-defined assumptions about degradation characteristics, which often fail to generalize to novel or compound corruptions Gao et al.
-(2024).
-(2) Prompt-based restoration methods, such as PromptIR Vaishnav et al.
-(2023) and DFPIR Tian et al.
-(2025a), rely on explicit, manually provided prompts to guide the restoration network by specifying degradation types.
-These methods offer flexible control but face a fundamental ”chicken-and-egg” dilemma: in real-world scenarios, the degradation type and location are rarely known beforehand, yet the network requires this information to perform effective restoration Jiang et al.
-(2025).
+(1) Adaptive feature learning methods, such as PromptIR (vaishnav2023promptir), ADAIR (cui2024adair) and AirNet (li2022all), utilize hand-crafted frequency priors or architectural inductive biases to automatically differentiate degradation types.
+Although these methods circumvent manual specification, they inherently depend on pre-defined assumptions about degradation characteristics, which often fail to generalize to novel or compound corruptions (gao2024prompt).
+(2) Prompt-based restoration methods, such as UniRestore (chen2025unirestore), InstructIR (conde2024high) and DFPIR (tian2025degradation), rely on explicit, manually provided prompts to guide the restoration network by specifying degradation types.
+These methods offer flexible control but face a fundamental ”chicken-and-egg” dilemma: in real-world scenarios, the degradation type and location are rarely known beforehand, yet the network requires this information to perform effective restoration (jiang2025survey).
 Moreover, TS instructions constrain generalization, particularly in cases of mixed degradations or varying homogeneous degradations (e.g., different noise levels).
@@
 To overcome these limitations, we reframe AIR as a latent prior inference problem.
-Unlike existing methods Vaishnav et al.
-(2023); Cui et al.
-(2024); Tian et al.
-(2025a) that rely on explicit degradation prompting or predefined architectural inductive biases, we propose learning degradation-aware representations directly from the degraded image through a multi-level feature descriptor inspired by the variational autoencoder (VAE) Kingma & Welling (2013).
+Unlike existing methods (chen2025unirestore; conde2024high; tian2025degradation) that rely on explicit degradation prompting or predefined architectural inductive biases, we propose learning degradation-aware representations directly from the degraded image through a multi-level feature descriptor inspired by the variational autoencoder (VAE) (kingma2013auto).
 Our learned prior eliminates the need for manual degradation hints and enables spatially adaptive restoration of diverse and unseen degradations.
@@
   • DAIR consistently outperforms SOTA on six common restoration tasks (e.g., snow, low-light), five compound degradations (e.g., haze + rain, low-light + haze + rain), and unseen degradations, achieving an average 1.68 dB PSNR gain.
-It also improves downstream tasks on images with unknown multi-type degradations, e.g., boosting YOLOv12-L Tian et al.
-(2025b) object detection by up to 7.40 mAP over SOTA AIR methods, highlighting strong generalizability (Fig.
+It also improves downstream tasks on images with unknown multi-type degradations, e.g., boosting YOLOv12-L (tian2025yolov12) object detection (OD) by up to 7.40 mAP over SOTA AIR methods, highlighting strong generalizability (Fig.
 1).
@@
 2.1 Image-to-image Translation
-Early image restoration methods were designed to address individual TS settings, such as denoising Zhang et al.
-(2017); Pang et al.
-(2021); Zhang et al.
-(2018a), dehazing Dong et al.
-(2020); Qin et al.
-(2020), derainingJiang et al.
-(2020); Ren et al.
-(2019), low-light image enhancement (LLIE) Wei et al.
-(2018); Yi et al.
-(2023), or deblurring Cho et al.
-(2021); Nah et al.
-(2017).
+Early image restoration methods were designed to address individual TS settings, such as denoising (zhang2017beyond; pang2021recorrupted; zhang2018ffdnet), dehazing (dong2020multi; qin2020ffa), deraining(jiang2020multi; ren2019progressive), low-light image enhancement (LLIE) (wei2018deep; yi2023diff), or deblurring (cho2021rethinking; nah2017deep).
 These task-dependent approaches achieved impressive results; their limited scope prevented generalization across diverse restoration challenges.
-Transformer-based models, such as Restormer Zamir et al.
-(2022) and Uformer Wang et al.
-(2022), have explored multiple degradation scenarios; however, they still require separate training for individual tasks.
+Transformer-based models, such as Restormer (zamir2022restormer) and Uformer (wang2022uformer), have explored multiple degradation scenarios; however, they still require separate training for individual tasks.
 To address this, recent methods shift towards AIR, aiming to handle diverse degradations within a unified framework.
-AirNet Li et al.
-(2022) pioneered this field by introducing contrastive learning to extract degradation representations, guiding restoration for unknown corruptions.
-ADAIR Cui et al.
-(2024) recalibrate features using adaptive frequency statistics, enabling task-agnostic restoration.
-Recent methods, such as DiffUIR Zheng et al.
-(2024), have also explored generative techniques (i.e., latent diffusion Rombach et al.
-(2022)).
+AirNet (li2022all) pioneered this field by introducing contrastive learning to extract degradation representations, guiding restoration for unknown corruptions.
+ADAIR (cui2024adair) recalibrate features using adaptive frequency statistics, enabling task-agnostic restoration.
+Recent methods, such as DiffUIR (zheng2024selective), have also explored generative techniques (i.e., latent diffusion (rombach2022ldm)).
 These unified approaches mark a significant step forward, minimizing TS training and advancing toward generalizable AIR solutions.
 2.2 Prompt-Guided All-in-one Restoration
-Image-to-image translation-based method commonly fails in separating unique degradations Brooks et al.
-(2023); Jiang et al.
-(2025).
-To counter this, recent AIR Gao et al.
-(2024); Vaishnav et al.
-(2023); Tian et al.
-(2025a) methods have incorporated MP to provide prior knowledge of degradation cues.
-PromptIR Vaishnav et al.
-(2023) integrates text-based prompts into a transformer framework, leveraging degradation-specific semantic information to guide the restoration process.
-OneRestore Guo et al.
-(2024) introduced visual and text prompts within a transformer-based framework to encode TS information.
-Recently, DFPIR Tian et al.
-(2025a) introduced a degradation-aware feature perturbation method that utilizes CLIP-encoded Radford et al.
-(2021) text prompts to guide channel shuffling and attention masking, enabling unified restoration across diverse degradations.
-UHD-Processor Liu et al.
-(2025) introduced a VAE-based framework with progressive frequency learning and MP for ultra-high-definition image restoration.
+Image-to-image translation-based method commonly fails in separating unique degradations (brooks2023instructpix2pix; jiang2025survey).
+To counter this, recent AIR (gao2024prompt; conde2024high; chen2025unirestore; tian2025degradation) methods have incorporated MP to provide prior knowledge of degradation cues.
+InstructIR (conde2024high) and UniRestore (chen2025unirestore) integrates text-based prompts into a transformer framework, leveraging degradation-specific semantic information to guide the restoration process.
+OneRestore (guo2024onerestore) introduced visual and text prompts within a transformer-based framework to encode TS information.
+Recently, DFPIR (tian2025degradation) introduced a degradation-aware feature perturbation method that utilizes CLIP-encoded (radford2021learning) text prompts to guide channel shuffling and attention masking, enabling unified restoration across diverse degradations.
+UHD-Processor (liu2025uhd) introduced a VAE-based framework with progressive frequency learning and MP for ultra-high-definition image restoration.
 Notably, unlike UHD-Processor and other prompt-based methods, our method learns latent degradation representations directly from the input, enabling blind restoration without user guidance.
-Figure 2: Overview of the proposed DAIR framework: (a) Overall architecture, (b) Latent prior encoding (which), (c) Degradation mapping (where), (d) Latent fusion (what), (e) 3WD restoration module.
-Best viewed with Zoom.
+Figure 2: Overview of the proposed DAIR framework.
+(a) The overall architecture consists of an encoder-decoder structure complemented by degradation priors from a VAE encoder.
+The degraded image is encoded by our Latent Prior Encoding (LPE) in both luminance and chrominance spaces.
+(b) Latent Prior Encoding (which) comprises multiple Degradation Aware Encoder Blocks (DAEB) that encode degradation prior information from the VAE.
+Latents encoded at different stages are passed to the Degradation Map blocks.
+(c) The Degradation Mapping (where) block utilizes LPE latents and VAE priors to generate degradation maps for each stage.
+(d) The Latent Fusion (what) block combines luminance and chrominance latents to generate unified latents.
 3 Method
@@
 Given a degraded input image $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$ with unknown degradation $\mathcal{D}_{R}$, our framework learn to infer stage-wise latent codes $x\ell$ ($\ell=0,\dots,3$) and a bottleneck global descriptor $\mu$, serving as degradation priors.
-The encoder employs a two-branch architecture that separately processes multi-scale structural features (e.g., luminance $L_{\ell}$) and color features (e.g., chrominance $C_{\ell}$), enabling more effective representation of both spatial and chromatic information Sharif et al.
-(2025); Yan et al.
-(2025).
+The encoder employs a two-branch architecture that separately processes multi-scale structural features (e.g., luminance $L_{\ell}$) and color features (e.g., chrominance $C_{\ell}$), enabling more effective representation of both spatial and chromatic information (sharif2025illuminatingdarknesslearningenhance; yan2025hvi).
 At each stage, $L_{\ell}$ and $C_{\ell}$ interact with $x_{\ell}$ via Degradation-Aware Encoder Blocks (DAEB), leveraging degradation information into the feature extraction process.
@@
 Concurrently, a mapping block transforms $(L_{\ell},C_{\ell},x_{\ell})$ into a compact DM $g_{\ell}$, which is also forwarded to the decoder for explicit degradation localization guidance.
-At the bottleneck, a $\mu$-guided fusion block merges the deepest LC features with the global degradation latent using adaptive scaling and shifting Perez et al.
-(2018), resulting in $F_{\mu}=\gamma\odot[L_{3}+C_{3}]+\beta$.
+At the bottleneck, a $\mu$-guided fusion block merges the deepest LC features with the global degradation latent using adaptive scaling and shifting (perez2018film), resulting in $F_{\mu}=\gamma\odot[L_{3}+C_{3}]+\beta$.
 The decoder reconstructs the clean image from $F_{\mu}$, leveraging the degradation focus map $g_{\ell}$ and the selected features $F_{\ell}$ derived from latent priors.
@@
 3.1 Learning Degradation Reasoning with Latent Prior
-We propose a multi-scale latent descriptor Kingma & Welling (2013) that directly infers continuous degradation representations from corrupted images without user-guided MP.
-Given $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$, the encoder extracts hierarchical feature maps $\{\mathbf{x}_{\ell}\}_{\ell=0}^{3}$ using convolutional layers and residual channel-attention Hu et al.
-(2018) blocks that emphasize degradation-relevant cues.
-At the deepest stage, multi-head self-attention Vaswani et al.
-(2017) refines the features to obtain the bottleneck representation $\mathbf{z}$.
+We propose a multi-scale latent descriptor (kingma2013auto) that directly infers continuous degradation representations from corrupted images without user-guided MP.
+Given $\mathbf{I}_{deg}\in\mathbb{R}^{3\times H\times W}$, the encoder extracts hierarchical feature maps $\{\mathbf{x}_{\ell}\}_{\ell=0}^{3}$ using convolutional layers and residual channel-attention (hu2018squeeze) blocks that emphasize degradation-relevant cues.
+At the deepest stage, multi-head self-attention (vaswani2017attention) refines the features to obtain the bottleneck representation $\mathbf{z}$.
 The VAE parameterizes a distribution with $\bm{\mu}=f_{\mu}(\mathbf{z})$ and $\log\bm{\sigma}^{2}=f_{\sigma}(\mathbf{z})$, sampling $\mathbf{z}_{reparam}=\bm{\mu}+\bm{\sigma}\odot\bm{\epsilon}$ where $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$.
@@
 3.1.1 Latent Prior Encoding
-Unlike existing methods (PromptIR Vaishnav et al.
-(2023), DFPIR Tian et al.
-(2025a), CAPTNet Gao et al.
-(2024)) that inject prompts during decoding, we inject VAE-derived degradation priors during encoding.
+Unlike existing methods (DFPIR (tian2025degradation), CAPTNet (gao2024prompt)) that inject prompts during decoding, we inject VAE-derived degradation priors during encoding.
 This enables our encoders to generate both degradation-aware and degradation-agnostic skip connections, providing richer guidance for reconstruction.
-We separately encode the luminance and chrominance (LC) components Smith (1978)($\mathbf{L}_{\text{deg}}$, $\mathbf{C}_{\text{deg}}$) of the degraded input using dedicated LC encoders Yan et al.
-(2025); Sharif et al.
-(2025) for better structural and chromatic information.
+We separately encode the luminance and chrominance (LC) components (smith1978color)($\mathbf{L}_{\text{deg}}$, $\mathbf{C}_{\text{deg}}$) of the degraded input using dedicated LC encoders (yan2025hvi; sharif2025illuminatingdarknesslearningenhance) for better structural and chromatic information.
 Fig.
@@
 3.1.2 Learnable Degradation Map
-Traditional skip connections propagate noise from degraded inputs to the decoder, without spatial awareness of the degradation Mao et al.
-(2016).
+Traditional skip connections propagate noise from degraded inputs to the decoder, without spatial awareness of the degradation (mao2016image).
 We counter these limitations with learnable DM that encode spatially-varying degradation patterns for effective reconstruction (Fig.
@@
 We first fuse complementary LC information: $\mathbf{f}^{LC}_{\ell}=\text{Conv}_{3\times 3}(L_{\ell}+C_{\ell})$.
-To capture frequency-domain degradation characteristics, we enhance features with spectral information via Fast Fourier Transform Cooley & Tukey (1965): $\mathcal{F}(L_{\ell})=\text{FFT}(L_{\ell}-\mu(L_{\ell}))$, $\mathcal{F}(C_{\ell})=\text{FFT}(C_{\ell}-\mu(C_{\ell}))$.
+To capture frequency-domain degradation characteristics, we enhance features with spectral information via Fast Fourier Transform (cooley1965algorithm): $\mathcal{F}(L_{\ell})=\text{FFT}(L_{\ell}-\mu(L_{\ell}))$, $\mathcal{F}(C_{\ell})=\text{FFT}(C_{\ell}-\mu(C_{\ell}))$.
 The magnitude and phase components are concatenated and projected:
@@
 As shown in Fig.
-2 (d), our fusion mechanism integrates three components: (i) intra-branch self-attention Sharif et al.
-(2025); Cai et al.
-(2023) for independent modality refinement, where $\hat{\mathbf{L}}_{3}=\text{MHSA}(\mathbf{L}_{3})$ and $\hat{\mathbf{C}}_{3}=\text{MHSA}(\mathbf{C}_{3})$; (ii) cross-branch attention Sharif et al.
-(2025) for complementary information exchange, where chrominance attends to luminance: $\mathbf{F}=\text{LayerNorm}\left(\text{MHCA}(\hat{\mathbf{C}}_{3},\hat{\mathbf{L}}_{3})+\hat{\mathbf{C}}_{3}+\hat{\mathbf{L}}_{3}\right)$, enabling the model to leverage structural and color information while preserving individual branch characteristics; and (iii) adaptive feature modulation conditioned on the global degradation prior $\mu$, combining LC features through adaptive scaling and shifting for TS feature modulation while maintaining underlying image characteristics.
-Unlike standard FiLM Perez et al.
-(2018) that applies uniform modulation, our approach makes degradation-aware reconstruction:
+2 (d), our fusion mechanism integrates three components: (i) intra-branch self-attention (sharif2025illuminatingdarknesslearningenhance; cai2023retinexformer) for independent modality refinement, where $\hat{\mathbf{L}}_{3}=\text{MHSA}(\mathbf{L}_{3})$ and $\hat{\mathbf{C}}_{3}=\text{MHSA}(\mathbf{C}_{3})$; (ii) cross-branch attention (sharif2025illuminatingdarknesslearningenhance) for complementary information exchange, where chrominance attends to luminance: $\mathbf{F}=\text{LayerNorm}\left(\text{MHCA}(\hat{\mathbf{C}}_{3},\hat{\mathbf{L}}_{3})+\hat{\mathbf{C}}_{3}+\hat{\mathbf{L}}_{3}\right)$, enabling the model to leverage structural and color information while preserving individual branch characteristics; and (iii) adaptive feature modulation conditioned on the global degradation prior $\mu$, combining LC features through adaptive scaling and shifting for TS feature modulation while maintaining underlying image characteristics.
+Unlike standard FiLM (perez2018film) that applies uniform modulation, our approach makes degradation-aware reconstruction:
 $[\bm{\gamma}_{\text{struct}},\bm{\beta}_{\text{color}}]$ $=\phi_{\mu}(\mu),\qquad\bm{\gamma}_{\text{struct}},\bm{\beta}_{\text{color}}\in\mathbb{R}^{B\times 320\times\frac{H}{8}\times\frac{W}{8}}$ (8)
@@
 $D_{\ell}=\phi_{\ell}(A_{\ell}\odot V_{\ell}+U_{\ell-1})$ (10)
-Notably, our proposed 3WD achieves linear computational complexity of $O(HW)$ per stage, in contrast to the quadratic complexity $O(H^{2}W^{2})$ of standard attention mechanisms Vaswani et al.
-(2017).
+Notably, our proposed 3WD achieves linear computational complexity of $O(HW)$ per stage, in contrast to the quadratic complexity $O(H^{2}W^{2})$ of standard attention mechanisms (vaswani2017attention).
 This efficiency stems from utilizing element-wise multiplication ($\odot$) between tensors of shape $H\times W\times C$, requiring exactly $HW\times C$ operations.
@@
 We evaluated our method under two degradation scenarios: common (non-overlapping) degradations and compound (overlapping) settings.
-For common degradations, we tested on standard tasks/datasets, including dehazing (SOTS Li et al.
-(2018)), deraining (Rain100L Fu et al.
-(2017)), desnowing (CCD Li et al.
-(2020)), deblurring Nah et al.
-(2017), deblurring (GoPro Nah et al.
-(2017)), denoising (DIV2K Agustsson & Timofte (2017)) with random noise levels ($\sigma\in[0,50]$) to improve generalization, and LLIE using the LSD dataset Sharif et al.
-(2025), captured in uncontrolled environments.
-In compound degradation, we employed the CDD dataset Guo et al.
-(2024), combining five degradations (e.g., haze+rain, low-light+haze+snow).
+For common degradations, we combined six widely used tasks, including dehazing (SOTS (li2018benchmarking)), deraining with heavy rain (Rain100H (fu2017deep)), desnowing (CCD (li2020desnownet)), real-world deblurring (nah2017deep), deblurring (GoPro (nah2017deep)), denoising (DIV2K (agustsson2017div2k)) with random noise levels ($\sigma\in[0,50]$) to improve generalization, and real-world LLIE using the LSD dataset (sharif2025illuminatingdarknesslearningenhance).
+In compound degradation, we employed the CDD dataset (guo2024onerestore), combining five degradations (e.g., haze+rain, low-light+haze+snow).
 Subsets like haze+snow and low-light+rain were reserved for testing generalization comparison on unseen degradations.
-Ablation and analysis, we combined four common tasks (denoising, desnowing, deraining, and LLIE) to analyze consistency.
-We compared DAIR against transformer-based baseline methods (Uformer Wang et al.
-(2022), Restormer Zamir et al.
-(2022)), SOTA AIR approaches (ADAIR Cui et al.
-(2024), AIRNet Li et al.
-(2022)), prompting-based methods (DFPIR Tian et al.
-(2025a), PromptIR Vaishnav et al.
-(2023)), and diffusion-based latent enhancement (DiffUIR Zheng et al.
-(2024)).
-Single-task benchmarks included deraining (HDCWNet Zhu et al.
-(2021), TransWeather Valanarasu et al.
-(2022)), Retinex-based LLIE (RetinexNet Wei et al.
-(2018), Diff-Retinex Yi et al.
-(2023)).
+We also included several real-world out-of-distribution datasets, such as SIDD for noisy images, LSD-U for unseen low-light scenarios, underwater image enhancement datasets, and medical image enhancement and denoising datasets.
+We compared DAIR against transformer-based baseline methods (Uformer (wang2022uformer), Restormer (zamir2022restormer)), SOTA AIR approaches (ADAIR (cui2024adair), AIRNet (li2022all)), prompting-based methods (DFPIR (tian2025degradation), PromptIR (vaishnav2023promptir)), and diffusion-based latent enhancement (DiffUIR (zheng2024selective)).
+Single-task benchmarks included deraining (HDCWNet (zhu2021hdcwnet), TransWeather (valanarasu2022transweather)), Retinex-based LLIE (RetinexNet (wei2018deep), Diff-Retinex (yi2023diff)).
 Implementation
-We first pre-train our latent encoder for 200,000 steps, combining six single and five compound (known) degradations using a composite loss (self-reconstruction + KL Kingma & Welling (2014) + discriminative latent regularizer Guo et al.
-(2023)).
+We first pre-train our latent encoder for 200,000 steps, combining six single and five compound (known) degradations using a composite loss (self-reconstruction + KL (kingma2014auto) + discriminative latent regularizer (guo2023letting)).
 The resulting encoder is frozen and reused for all single, compound, and unseen degradation experiments without any fine-tuning.
@@
 4.2.1 Multi-task (6D) Restoration
-Table 1: Performance comparison across different image restoration tasks. Best results in bold red, second best underlined, and increment over best performing method highlighted in blue.
+Table 1: Performance comparison across six different image restoration tasks under all-in-one setting: a unified model is trained on a combined set of images obtained from all degradation types and levels.
+Best results in bold red, second best underlined, and increment over best performing method highlighted in blue.
 Method MP Params (M) GFLOPs [^1] Computed on $256\times 256$ images. Lowlight Dehazing Denoising Desnowing Deblurring Deraining Average
-PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
-Uformer Wang et al. (2022) 20.63 43.86 12.10/0.4897 29.15/0.9727 27.55/0.8201 24.80/0.8832 23.67/0.8074 28.74/0.8823 24.34/0.8092
-Restormer Zamir et al. (2022) 26.13 141.75 15.33/0.6180 28.51/0.9738 20.41/0.5357 22.98/0.8585 25.41/0.8286 25.99/0.8164 23.10/0.7718
-AIRNet Li et al. (2022) 8.93 301.27 8.87/0.3655 25.92/0.9663 28.87/0.8789 19.08/0.7300 22.81/0.8067 24.92/0.8027 21.75/0.7583
-PromptIR Vaishnav et al. (2023) 35.59 158.14 14.73/0.5374 30.41/0.9580 23.96/0.6378 27.08/0.9169 26.04/0.8366 30.01/0.8992 25.37/0.7977
-DiffUIR Zheng et al. (2024) 36.26 4400.00 10.76/0.4101 30.05/0.9797 20.25/0.6859 20.39/0.8014 25.03/0.8474 27.46/0.8917 22.32/0.7694
-ADAIR Cui et al. (2024) 28.78 147.18 12.76/0.4974 30.15/0.9783 28.65/0.8715 25.86/0.8812 25.14/0.8200 29.56/0.8901 25.35/0.8231
-DFPIR Tian et al. (2025a) 31.07 151.07 14.84/0.5232 28.29/0.9709 27.78/0.8341 21.74/0.7935 26.10/0.8271 28.02/0.8372 24.46/0.7977
-DAIR (Ours) 18.08 45.65 16.87/0.6661 34.08/0.9864 29.12/0.8971 31.50/0.9570 26.82/0.8793 30.51/0.9117 28.15/0.8829
-Improvement -10.70 -96.10 +1.54/+0.0481 +3.67/+0.0067 +0.25/+0.0182 +4.42/+0.0401 +0.72/+0.0427 +0.50/+0.0125 +2.78/+0.0598
-Figure 3: Visual comparison for six common degradation settings. The proposed method produces consistent and plausible images compared to the existing methods.
-Table 1 and Figure 3 illustrate DAIR’s performance across six common degradations.
+PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓
+Uformer 20.63 43.86 12.10/0.4897/0.3621 29.15/0.9727/0.0195 27.55/0.8201/0.1334 24.80/0.8832/0.0904 23.67/0.8074/0.2401 28.74/0.8823/0.1051 24.34/0.8092/0.1584
+Restormer 26.13 141.75 15.33/0.6180/0.2697 28.51/0.9738/0.0154 20.41/0.5357/0.4148 22.98/0.8585/0.1213 25.41/0.8286/0.2080 25.99/0.8164/0.1547 23.10/0.7718/0.1973
+AIRNet 8.93 301.27 8.87/0.3655/0.3679 25.92/0.9663/0.0237 28.87/0.8789/0.1020 19.08/0.7300/0.2596 22.81/0.8067/0.2284 24.92/0.8027/0.1692 21.75/0.7583/0.1918
+PromptIR 35.59 158.14 14.73/0.5374/0.3665 30.41/0.9580/0.0221 23.96/0.6378/0.2339 27.08/0.9169/0.0634 26.04/0.8366/0.2069 30.01/0.8992/0.0669 25.37/0.7977/0.1600
+DiffUIR 36.26 4400.00 10.76/0.4101/0.4290 30.05/0.9797/0.0151 20.25/0.6859/0.2329 20.39/0.8014/0.1709 25.03/0.8474/0.2019 27.46/0.8917/0.0719 22.32/0.7694/0.1870
+ADAIR 28.78 147.18 12.76/0.4974/0.3605 30.15/0.9783/0.0135 28.65/0.8715/0.1043 25.86/0.8812/0.0966 25.14/0.8200/0.2307 29.56/0.8901/0.0889 25.35/0.8231/0.1491
+DFPIR 31.07 151.07 14.84/0.5232/0.3766 28.29/0.9709/0.0214 27.78/0.8341/0.1433 21.74/0.7935/0.2189 26.10/0.8271/0.2230 28.02/0.8372/0.1506 24.46/0.7977/0.1890
+DAIR (Ours) 18.08 45.65 16.87/0.6661/0.2559 34.08/0.9864/0.0100 29.12/0.8971/0.0826 31.50/0.9570/0.0264 26.82/0.8793/0.1653 30.51/0.9117/0.0663 28.15/0.8829/0.1011
+Improvement -10.70 -96.10 +1.54/+0.0481/-0.0138 +3.67/+0.0067/-0.0035 +0.25/+0.0182/-0.0194 +4.42/+0.0401/-0.0370 +0.72/+0.0427/-0.0366 +0.50/+0.0125/-0.0006 +2.78/+0.0598/-0.0480
+Figure 3: Visual comparison for six common degradation settings under the all-in-one setting. The proposed method produces consistent and plausible images compared to the existing methods.
+Table 1 and Figure 3 illustrate DAIR’s performance across six common degradations under all-in-one setting.
 Achieving the highest average PSNR (28.15) and SSIM (0.8829), DAIR delivers significant improvements, including +5.64 PSNR and +0.0758 SSIM in desnowing, and +1.54 PSNR and +0.0481 SSIM in lowlight enhancement.
-Despite requiring only 45G FLOPs, DAIR outperforms computationally intensive methods like ADAIR Cui et al.
-(2024)(147G FLOPs) while operating without MP, ensuring scalability and adaptability.
+Despite requiring only 45G FLOPs, DAIR outperforms computationally intensive methods like ADAIR (cui2024adair)(147G FLOPs) while operating without MP, ensuring scalability and adaptability.
 By leveraging degradation-aware latent priors, DAIR generates robust, high-quality restorations across diverse degradation types, making it ideal for real-world applications like autonomous systems and healthcare imaging.
 4.2.2 Single-task Evaluation: LLIE and Desnowing
-Table 2: Performance comparison on desnowing and LLIE tasks. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
+Table 2: Performance comparison on desnowing and LLIE tasks.
 (a) Desnowing (b) Low-light enhancement
-Method PSNR↑/SSIM↑ Method PSNR↑/SSIM↑
-HDCWNet Zhu et al. (2021) 29.25/0.9171 RetinexNet Wei et al. (2018) 14.19/0.5183
-TransWeather Valanarasu et al. (2022) 23.30/0.7631 Diff-Retinex Yi et al. (2023) 15.38/0.5038
-Uformer Wang et al. (2022) 28.86/0.9324 Uformer Wang et al. (2022) 13.05/0.5716
-Restormer Zamir et al. (2022) 23.21/0.8686 Restormer Zamir et al. (2022) 15.78/0.6184
-ADAIR Cui et al. (2024) 28.31/0.9300 ADAIR Cui et al. (2024) 15.19/0.6094
-DAIR (Ours) 32.75/0.9632 DAIR (Ours) 16.30/0.6634
-Improvement +3.50/+0.0308 Improvement +0.52/+0.0450
+Method PSNR↑/SSIM↑/LPIPS↓ Method PSNR↑/SSIM↑/LPIPS↓
+HDCWNet 29.25/0.9171/0.0517 RetinexNet 14.19/0.5183/0.3812
+TransWeather 23.30/0.7631/0.1739 Diff-Retinex 15.38/0.5038/0.3841
+Uformer 28.86/0.9324/0.0904 Uformer 13.05/0.5716/0.3621
+Restormer 23.21/0.8686/0.1059 Restormer 15.78/0.6184/0.3372
+ADAIR 28.31/0.9300/0.0457 ADAIR 15.19/0.6094/0.3325
+DAIR (Ours) 32.75/0.9632/0.0213 DAIR (Ours) 16.30/0.6634/0.3279
+Improvement +3.50/+0.0308/-0.0244 Improvement +0.52/+0.0450/-0.0046
 We evaluated DAIR on single restoration tasks to demonstrate its adaptability.
-Table 2 shows DAIR achieves 32.75 PSNR / 0.9632 SSIM for desnowing, outperforming HDCWNet Zhu et al.
-(2021) by +3.50 PSNR / +0.0308 SSIM.
-For LLIE, DAIR achieves 16.30 PSNR / 0.6634 SSIM, improving upon Restormer Zamir et al.
-(2022) by +0.52 PSNR / +0.0450 SSIM.
+Table 2 shows DAIR achieves 32.75 PSNR / 0.9632 SSIM for desnowing, outperforming HDCWNet (zhu2021hdcwnet) by +3.50 PSNR / +0.0308 SSIM.
+For LLIE, DAIR achieves 16.30 PSNR / 0.6634 SSIM, improving upon Restormer (zamir2022restormer) by +0.52 PSNR / +0.0450 SSIM.
 By leveraging degradation-aware latent priors, DAIR can also adapt to varying degradation types without external MP.
 4.3 Results on Compound Degradation
-Table 3: Performance comparison across five compound degradation. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
+Table 3: Quantitative evaluation of restoration performance on five compound degradation types, derived from combinations of low-light, rain, haze, and snow datasets.
 Method Haze + Rain Low + Haze Low + Haze + Rain Low + Haze + Snow Low + Snow Average
-PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
-Uformer Wang et al. (2022) 22.49/0.8889 21.37/0.8252 20.61/0.7838 18.64/0.6802 20.31/0.7150 20.68/0.7786
-Restormer Zamir et al. (2022) 23.57/0.8825 21.44/0.7753 20.53/0.7265 18.46/0.6494 19.97/0.6664 20.79/0.7400
-AIRNet Li et al. (2022) 18.58/0.7977 11.73/0.6905 13.58/0.6320 12.87/0.5898 13.76/0.5759 14.10/0.6572
-PromptIR Vaishnav et al. (2023) 21.78/0.8818 20.63/0.8231 19.83/0.7705 18.42/0.6801 20.74/0.7070 20.28/0.7725
-DiffUIR Zheng et al. (2024) 18.42/0.8022 14.28/0.7150 15.79/0.6568 15.33/0.6238 12.60/0.6292 15.28/0.6854
-ADAIR Cui et al. (2024) 23.51/0.8966 22.78/0.8420 21.06/0.7908 20.46/0.7229 21.35/0.7215 21.83/0.7948
-DFPIR Tian et al. (2025a) 20.81/0.8308 19.01/0.7784 18.82/0.7139 15.58/0.6307 19.70/0.6402 18.78/0.7188
-DAIR (Ours) 25.25/0.9259 23.15/0.8541 22.03/0.8200 20.81/0.7613 21.79/0.7772 22.61/0.8277
-Improvement +1.68/+0.0293 +0.37/+0.0121 +0.97/+0.0292 +0.35/+0.0384 +0.44/+0.0557 +0.78/+0.0329
+PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓
+Uformer 22.49/0.8889/0.0867 21.37/0.8252/0.1695 20.61/0.7838/0.1908 18.64/0.6802/0.3052 20.31/0.7150/0.2513 20.68/0.7786/0.2007
+Restormer 23.57/0.8825/0.0959 21.44/0.7753/0.1884 20.53/0.7265/0.2367 18.46/0.6494/0.3309 19.97/0.6664/0.3003 20.79/0.7400/0.2304
+AIRNet 18.58/0.7977/0.1987 11.73/0.6905/0.2587 13.58/0.6320/0.3731 12.87/0.5898/0.4141 13.76/0.5759/0.4372 14.10/0.6572/0.3364
+PromptIR 21.78/0.8818/0.1012 20.63/0.8231/0.1850 19.83/0.7705/0.2223 18.42/0.6801/0.3171 20.74/0.7070/0.2645 20.28/0.7725/0.2180
+DiffUIR 18.42/0.8022/0.1758 14.28/0.7150/0.2562 15.79/0.6568/0.3308 15.33/0.6238/0.3500 12.60/0.6292/0.2988 15.28/0.6854/0.2823
+ADAIR 23.51/0.8966/0.0791 22.78/0.8420/0.1419 21.06/0.7908/0.1774 20.46/0.7229/0.2416 21.35/0.7215/0.2444 21.83/0.7948/0.1769
+DFPIR 20.81/0.8308/0.1548 19.01/0.7784/0.2339 18.82/0.7139/0.2800 15.58/0.6307/0.3366 19.70/0.6402/0.3227 18.78/0.7188/0.2656
+DAIR (Ours) 25.25/0.9259/0.0637 23.15/0.8541/0.1408 22.03/0.8200/0.1618 20.81/0.7613/0.2233 21.79/0.7772/0.1874 22.61/0.8277/0.1554
+Improvement +1.68/+0.0293/-0.0154 +0.37/+0.0121/-0.0011 +0.97/+0.0292/-0.0156 +0.35/+0.0384/-0.0183 +0.44/+0.0557/-0.0570 +0.78/+0.0329/-0.0215
 Figure 4: Visual comparison for five compound degradations. The proposed DAIR can handle compound degradation and produce visually pleasing images.
-We evaluated DAIR on compound degradation scenarios.
-As shown in Table 3, DAIR achieves the highest average PSNR of 22.61 and SSIM of 0.8277, outperforming the next-best method (ADAIR Cui et al.
-(2024)) by +0.78 PSNR / +0.0329 SSIM.
+We evaluated DAIR on five compound degradation scenarios derived from combinations of low-light, rain, haze, and snow datasets.
+As shown in Table 3, DAIR achieves the highest average PSNR of 22.61 and SSIM of 0.8277, outperforming the next-best method (ADAIR (cui2024adair)) by +0.78 PSNR / +0.0329 SSIM.
 Notable improvements include ”Haze + Rain” (+1.68 PSNR / +0.0293 SSIM) and ”Low + Snow” (+0.44 PSNR / +0.0557 SSIM).
@@
 4.4 Latent prior Guided Unseen Task Restoration
-Table 4: Performance comparison on unknown compound degradations. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
+Our latent prior descriptor enables robust handling of unknown degradations while maintaining clear separation from known ones.
+As shown in Table 5, the VAE-based encoder learns degradation characteristics directly from corrupted images, achieving strong clustering for both seen and unseen degradations with KNN accuracy of 0.994/0.976, separation ratios of 2.911/1.523, and optimal global metrics (lowest Davies-Bouldin index, highest Calinski-Harabasz score).
+This learned prior guides the main network with degradation-specific descriptions, enabling clear latent separation for different degradation, where existing methods fail (Fig.
+5(b)).
+Consequently, DAIR restores unknown compound degradations effectively (Fig.
+5(a)), achieving average gains of +1.47 dB PSNR and +0.0412 SSIM over SOTA methods (Table 5).
+Please refer appendix for more details.
+Table 4: Latent prior descriptor yields strong clustering and separation for seen/unseen degradations.
+Model Seen Unseen Overall Global Metrics
+Acc↑ Sep↑ Acc↑ Sep↑ Acc↑ Sep↑ DB↓ CH↑
+w/o SupCon 0.864 2.00 0.787 1.32 0.874 1.79 2.69 254.48
+Steps = 0 0.826 3.11 0.756 1.33 0.844 2.28 4.84 458.34
+Steps = 100k 0.871 2.36 0.828 1.27 0.892 1.95 2.45 289.84
+Full Training 0.994 2.91 0.976 1.52 0.991 3.22 1.23 549.34
+Table 5: Performance comparison on unknown compound degradations.
 Method Haze + Snow Low + Rain Average
 PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
-ADAIR Cui et al. (2024) 16.07/0.8130 21.67/0.7877 18.87/0.8003
-PromptIR Vaishnav et al. (2023) 17.84/0.8032 21.15/0.7752 19.50/0.7892
-DFPIR Tian et al. (2025a) 14.50/0.7577 21.52/0.7021 18.01/0.7299
+ADAIR 16.07/0.8130 21.67/0.7877 18.87/0.8003
+PromptIR 17.84/0.8032 21.15/0.7752 19.50/0.7892
+DFPIR 14.50/0.7577 21.52/0.7021 18.01/0.7299
 DAIR (Ours) 18.65/0.8531 23.29/0.8299 20.97/0.8415
 Improvement +0.81/+0.0499 +1.62/+0.0422 +1.47/+0.0523
-The primary motivation for leveraging latent priors is to address unknown degradations while maintaining a clear distinction from known degradations.
-Fig.
-5 highlights the restoration of unknown compound degradations, comparing DAIR with SOTA methods.
-Our approach effectively handles these degradations, as demonstrated by t-SNE visualization in Fig.
-5 (b).
-Table 4 summarizes the results, showing that our method achieves notable improvements over existing approaches, with average gains of +1.47 dB PSNR and +0.0412 SSIM for unseen degradations.
 Refer to caption Refer to caption Refer to caption Refer to caption (b)
 Figure 5: DAIR performance on unseen tasks: (a) Visual results for unseen compound degradation (low-light + rain); (b) t-SNE embeddings showing separation of unseen and known degradations.
-4.5 Degradation Map
-Figure 6: Visualization and restoration with the proposed DM.
-Our proposed DM effectively learns spatially correlated regions, enabling the decoder to focus on ”where” to reconstruct the degraded image.
+4.5 Analysis of Reasoning Restoration
+4.5.1 Latent Prior Encoding
+The distance matrix in Fig.
+7 (a) shows aggregated latent representations: strong intra-class consistency (diagonal: 160–194) and clear inter-class separation.
+Related degradations have moderate distances (e.g., Haze–Lowlight: 102, Haze–Blur: 86), while dissimilar pairs show larger separations (e.g., Snow–Lowlight: 201, Snow–Haze: 136).
 As shown in Fig.
-6, the comparison with traditional skip connections highlights that our learnable approach precisely identifies affected areas and significantly outperforms conventional methods.
-4.6 Restoration with 3WD
-We evaluate DAIR with 3WD against self-attention (SA) Vaswani et al.
-(2017) and cross-attention (CA) Vaswani et al.
-(2017) baselines.
-As shown in Table 5, DAIR with 3WD consistently outperforms both baselines, achieving average improvements of +7.72 dB PSNR and +0.2605 SSIM, with significant gains in desnowing (+11.89 dB) and denoising (+8.24 dB).
-Additionally, DAIR operates with linear complexity $O(d^{2}\times H\times W)$ compared to the baselines’ quadratic $O(d\times H^{2}\times W^{2})$, delivering 85-257× speedup and 76-89% memory reduction while answering our reasoning cues.
-Table 5: Performance comparison of proposed 3WD restoration with existing attention mechanism.
-Best results in bold red, second best underlined.
-Improvement over the best performing method highlighted in blue.
-Method Complexity Denoising Desnowing Deraining Lowlight Average
-(Speedup) PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
-DAIR + SA Vaswani et al. (2017) $O(H^{2}W^{2})$ 18.90/0.6213 19.70/0.6738 22.94/0.6541 15.16/0.4348 19.18/0.5960
-DAIR + CA Vaswani et al. (2017) $O(H^{2}W^{2})$ 20.87/0.6290 16.07/0.6469 20.90/0.6448 13.00/0.3854 17.71/0.5765
-DAIR + 3WD (Ours) $O(HW)$ 29.11/0.8971 31.59/0.9581 30.21/0.9070 16.68/0.6638 26.90/0.8565
-Improvement Linear scaling +8.24/+0.2681 +11.89/+0.2843 +7.27/+0.2529 +1.52/+0.2290 +7.72/+0.2605
+7 (b), encoder stages learn progressive feature selection based on these latent priors: early stages (L1–L2) show higher distances ( 180–200) indicating stronger luminance preference for structural information, while deeper stages (L3–L4) show lower distances ( 60–100) reflecting increased chrominance preference for color-specific restoration.
+This degradation-aware hierarchical guidance enables DAIR to adaptively leverage structural and chromatic information across different encoder depths for different degradations, validating our ”which” reasoning.
+Figure 6: Latent prior encoding validation. (a) Degradation separability via distance matrix. (b) Encoder progression: structural (L1–L2) to chromatic (L3–L4) features.
+Figure 7: Degradation localization. Our learnable map precisely identifies affected regions (yellow box).
+4.5.2 Degradation Map
+Table 6 quantitatively evaluates our degradation map localization accuracy (”where”).
+To evaluate degradation localization, we generate ground truth masks by computing pixel-wise absolute differences $|I_{\text{degraded}}-I_{\text{clean}}|$ and compare them with our predicted degradation maps.
+Our method achieves a mean IoU of 0.861±0.169 and a Dice score of 0.912±0.118.
+Pixel-level degradations (Noise, Lowlight, Haze) achieve near-perfect localization (IoU >> 0.95), while spatially-varying degradations (Rain, Snow, Blur) maintain strong performance despite higher variance.
+As shown in Fig.
+7, our learnable degradation maps precisely identify affected regions and significantly outperform traditional skip connections, enabling the decoder to focus reconstruction on degraded areas.
+Table 6: Quantitative evaluation of degradation map localization (”where”) using IoU and Dice metrics across pixel-wise and spatially-varying degradations.
+Metric Noise Lowlight Haze Rain Snow Blur Mean ± Std
+[…]
+[…]
+4.5.3 Latent Encoded Fusion
+Table 7: Latent Prior Encoding Quality: Quantitative evaluation of degradation-aware representation learning across different degradation types
+Metric Blur Haze Lowlight Noise Rain Snow GLOBAL
+[…]
+[…]
+Table 7 validates our latent fusion for degradation semantic encoding (”what”).
+By fusing luminance and chrominance representations with global priors via adaptive modulation, our method achieves a separation ratio of 1.56 and KNN accuracy of 0.932.
+Frequency-domain degradations (blur, noise) show perfect classification (Acc = 1.0), while pixel-level and spatially-varying degradations maintain strong separability (Acc >> 0.88, Sep >> 1.15).
+These results demonstrate that our latent fusion captures degradation-specific characteristics without explicit prompts, guiding the decoder in determining ”what” content to reconstruct.
+4.5.4 Restoration with 3WD
+Table 8: Comparison and restoration with 3WD.
+Method Complexity Memory (MB)↓ FPS↑ PSNR↑/SSIM↑
+CA $O(H^{2}W^{2})$ - - 17.71/0.5765
+SA $O(H^{2}W^{2})$ - - 19.05/0.5444
+WA $O(M^{2}\cdot HW)$ 4519.90 8.77 23.42/0.7944
+3WD (Ours) $O(HW)$ 2045.94 45.69 26.90/0.8565
+Improvement Linear -54.7% +5.2× +3.48/+0.0621
+We evaluate DAIR with 3WD for leveraging reasoning cues against attention-based baselines: self-attention (SA), cross-attention (CA) (vaswani2017attention), and window attention (WA) (liang2021swinir).
+As shown in Table 8, our method achieves 26.90 dB PSNR and 0.8565 SSIM, outperforming the best baseline (WA) by +3.48 dB and +0.0621 SSIM.
+Critically, our linear complexity $O(HW)$ enables 5.2× speedup and 54.7% memory reduction compared to WA’s quadratic $O(M^{2}\cdot HW)$ complexity.
+We achieve 45.69 FPS with 4593.76 MB at HD 720p, with similar efficiency maintained at 2K resolution (20.45 FPS, 4573.29 MB).
+This validates that our 3WD reasoning mechanism effectively captures spatial-degradation dependencies without quadratic computational overhead.
 4.6 Ablation Study
-We conducted an ablation study to evaluate DAIR’s key components: Latent Priors Encoding (LPE) (”which”), Degradation Map (DM) (”where”), Latent fusion (LF) (”what”), and Restoration Block (3WD).
+Table 9: Ablation study on key components. Tick () indicates the component is used, cross () indicates not used.
+Method LP LF DM 3WD PSNR↑/SSIM↑
+Base Model 20.22/0.6936
+Base + LP 20.49/0.7025
+Base + LP + LF 21.08/0.7208
+Base + LP + LF + DM 21.77/0.7439
+DAIR w/o Latent Prior 22.67/0.7689
+DAIR 26.90/0.8565
+Improvement +6.68/+0.1629
+We conducted an ablation study combining four challenging tasks (denoising, desnowing, deraining, LLIE) to evaluate the key components of DAIR: Latent Priors Encoding (LPE) (”which”), Degradation Map (DM) (”where”), Latent Fusion (LF) (”what”), and Restoration Block (3WD).
 As shown in Table 9, the inclusion of the proposed component progressively enhances performance, with DAIR achieving the highest average PSNR of 26.90 and SSIM of 0.8565.
 Notably, removing Latent Priors results in significantly worse performance (-4.23 PSNR / -0.0876 SSIM), highlighting its critical role in capturing detailed degradation descriptions, subsequently guiding the model in effectively segregating different types of degradations.
-Table 6: Ablation study on key components. Tick () indicates the component is used, cross () indicates not used. Best results in bold red.
-Method LP LF DM 3WD Denoise Desnowing Derain Lowlight Average
-PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
-Base Model 23.42/0.7237 20.76/0.7696 22.70/0.7698 13.99/0.5111 20.22/0.6936
-Base + LP 24.50/0.7417 20.75/0.7783 22.46/0.7658 14.23/0.5240 20.49/0.7025
-Base + LP + LF 25.83/0.7799 21.21/0.8014 22.23/0.7545 15.04/0.5474 21.08/0.7208
-Base + LP + LF + DM 26.80/0.8158 21.97/0.8442 22.79/0.7614 15.52/0.5540 21.77/0.7439
-DAIR w/o Latent Prior 26.07/0.8855 26.83/0.8306 22.76/0.7741 15.00/0.5853 22.67/0.7689
-DAIR 29.11/0.8971 31.59/0.9581 30.21/0.9070 16.68/0.6638 26.90/0.8565
-Improvement +5.69/+0.1734 +10.83/+0.1885 +7.51/+0.1372 +2.69/+0.1527 +6.68/+0.1629
+Please see Appendix B for module-wise ablation and details.
+4.7 Real-world Implications
+Table 10: Object Detec-
+tion on restored images.
+Method Average Precision
+AP50:95 AP50 AP75
+PromptIR 30.9 33.3 31.9
+DFPIR 27.9 30.0 29.3
+ADAIR 28.9 31.2 30.2
+DAIR (Ours) 34.9 37.3 36.4
+Improvement +4.0 +4.0 +4.5
+Table 11: Real-world unseen restoration.
+Method NIQE↓/MUSIQ↑/BRISQUE↓
+PromptIR 9.65/43.81/40.03
+DFPIR 8.01/45.02/25.52
+ADAIR 6.85/46.10/28.45
+DAIR (Ours) 5.25/50.19/25.97
+Improvement -1.60/+4.09/+0.45
+Tables 11 and 11 evaluate DAIR’s restoration quality through downstream tasks and real-world unseen restoration.
+For OD (Table 11), YOLOv12 tested on degraded MS-COCO images shows DAIR achieves 34.9 AP50:95, 37.3 AP50, and 36.4 AP75, outperforming PromptIR by +4.0, +4.0, and +4.5 points, demonstrating preserved semantic content.
+For perceptual quality (Table 11), no-reference metrics across real-world unseen datasets show DAIR achieves best scores: 5.25 NIQE (-1.60), 50.19 MUSIQ (+4.09), and 25.97 BRISQUE (+0.45), confirming superior perceptual quality and strong generalization to unseen out-of-distribution scenarios.
+See Appendix D for details.
 5 Conclusion
@@
 Architecture.
-To effectively integrate the VAE into our framework, we design a hybrid U-Net encoder–decoder with multi-head self-attention Tian et al.
-(2025a) at the bottleneck.
+To effectively integrate the VAE into our framework, we design a hybrid U-Net encoder–decoder with multi-head self-attention (tian2025degradation) at the bottleneck.
 In particular, both the hybrid VAE and the reconstruction network incorporate the same spatial feature dimensions, enabling seamless interaction between latent representations and restoration features.
 The encoder progressively downsamples RGB images through four stages with channel dimensions $[40,80,160,320]$ using 3×3 convolutions (stride=2 for downsampling, padding=1).
-Each stage contains a Residual Attention Block He et al.
-(2016) combining spatial convolutions with channel attention Hu et al.
-(2018) (reduction ratio 8):
+Each stage contains a Residual Attention Block (he2016resnet) combining spatial convolutions with channel attention (hu2018squeeze) (reduction ratio 8):
 $\text{ResAttn}(\mathbf{f})=\mathbf{f}+\text{Conv}_{3\times 3}(\text{ReLU}(\text{Conv}_{3\times 3}(\mathbf{f})))\odot\sigma(\text{Conv}_{1\times 1}(\text{GAP}(\mathbf{f})))$ (11)
@@
 where $\mathbf{I}_{deg}$ is the input degraded image and $\hat{\mathbf{I}}_{deg}=D_{\theta}(\mathbf{z})$ is the reconstructed output from the decoder.
-The KL divergence Kullback & Leibler (1951) term regularizes the latent distribution:
+The KL divergence (kullback1951information) term regularizes the latent distribution:
 $\mathcal{L}_{\text{KL}}=\text{KL}(q_{\phi}(\mathbf{z}|\mathbf{I}_{deg})\|\mathcal{N}(\mathbf{0},\mathbf{I}))$ (16)
 This loss term encourages the learned latent distribution $q_{\phi}(\mathbf{z}|\mathbf{I}_{deg})$ to be close to a standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, ensuring that the latent space has good sampling properties and prevents overfitting.
-To encourage discriminative latent representations, we incorporate weakly Supervised Contrastive Learning Guo et al. (2023) on the spatially-pooled encoder:
+To encourage discriminative latent representations, we incorporate weakly Supervised Contrastive Learning (guo2023letting) on the spatially-pooled encoder:
 $\mathcal{L}_{\text{SupCon}}=-\sum_{i}\log\frac{\sum_{j\in P_{i}}\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{j}/\tau)}{\sum_{k\neq i}\exp(\mathbf{z}_{i}\cdot\mathbf{z}_{k}/\tau)}$ (17)
@@
 This loss pulls together samples from the same degradation category while pushing apart samples from different categories in the latent space, encouraging the encoder to learn discriminative representations for different types of image degradations.
+Capture degradation with VAE.
+Fig.
+8 illustrates the proposed VAE’s progressive convergence and degradation-aware representation learning (Step = 0 → 100k).
+Incorporating SupCon loss as weak supervision improves latent separability of corruption types while preserving reconstruction fidelity, enabling the model to disentangle mixed degradations and generalize beyond the training distribution.
+Notably, it separates the unseen “Low + Rain” (guo2024onerestore) from the related seen “Low + Haze + Rain,” and characterizes out-of-domain cases such as underwater images, evidencing robust clustering and stable training dynamics.
+Figure 8: Learning degradation with VAE. (a) VAE without SupCon. (b) VAE at Step = 0. (c) VAE at Step = 100k. (d) VAE (fully trained).
 A.3 DAIR Details
@@
 22:return $\theta_{\text{VAE}},\theta_{\text{REST}}$
-Appendix B Additional Experimental Results
-B.1 Quantitative Comparison
-B.1.1 Low-light
-We utilize the LSD dataset Sharif et al.
-(2025), collected in uncontrolled low-light settings, offering diverse indoor and outdoor scenes under varying conditions.
+Appendix B Analysis and Ablation of DAIR
+We conduct an ablation and module-wise analysis on four common tasks: denoising, desnowing, deraining, and LLIE.
+This section provides a detail the performance and contributions of our proposed module across these tasks.
+B.1 Latent Prior Encoding (”which”)
+We analyze the “which” component and demonstrate that LC separation makes feature selection both explicit and effective.
+By routing structure-dominant cues through luminance and color-specific cues through chrominance, the encoder preserves complementary statistics that can be modulated by latent priors per branch.
+This design validates our reasoning by enabling degradation-aware, stage-wise selection of “which” features to leverage, resulting in substantial gains over luminance-only, chrominance-only, and RGB based encoders, as shown in Table 12.
+Table 12: Ablation study of which module.
+Method PSNR↑/SSIM↑/LPIPS↓
+Luma only 23.58/0.7996/0.1946
+Chroma only 23.41/0.7927/0.2014
+RGB 23.44/0.8002/0.1892
+Proposed (Both) 26.90/0.8565/0.1316
+Improvement +3.32/+0.0569/-0.0630
+B.2 Degradation Map (”where”)
+We evaluate the impact of the degradation map (”where”) separately.
+Table 13 demonstrates that the proposed degradation map with the FFT module significantly improves restoration performance.
+The model without the degradation map achieved a PSNR of 22.80, SSIM of 0.7769, and LPIPS of 0.2116, whereas our approach with the degradation map improved these metrics to a PSNR of 26.90, SSIM of 0.8565, and LPIPS of 0.1316, resulting in enhancements of +4.10 in PSNR, +0.0796 in SSIM, and a reduction of -0.0800 in LPIPS.
+It also ensures that the degradation map captures frequency-domain degradation characteristics, leveraging frequency cues.
+Table 13: Ablation study of degradation map (”where”) module.
+Method PSNR↑/SSIM↑/LPIPS↓
+w/o FFT 22.80/0.7769/0.2116
+Proposed 26.90/0.8565/0.1316
+Improvement +4.10/+0.0796/-0.0800
+B.3 Latent Fusion (”what”)
+We systematically evaluate the impact of different fusion strategies on restoration performance.
+Table 14 illustrates the results of latent fusion across various settings.
+Our full $\mu$-fusion (Factor=1) achieves a PSNR of 26.90 dB, outperforming the second-best method by +3.92 dB.
+Notably, removing degradation-aware components (beta: 22.97 dB, MU-fusion: 22.74 dB) or employing fixed fusion factors (Factor=0: 17.88 dB) results in significant performance drops.
+This underscores that the learned $\mu$-fusion weights effectively capture degradation characteristics and facilitate semantically meaningful decisions.
+Table 14: Ablation study of fusion strategies. Best in bold red, second underlined.
+Method Beta Alpha Factor PSNR↑/SSIM↑/LPIPS↓
+w/o MU 22.74/0.7791/0.2135
+w/o beta 1 22.97/0.7824/0.2147
+alpha + beta 0 17.88/0.7272/0.2488
+Fusion 0.1 0.1 16.28/0.7350/0.2345
+Fusion 0.5 0.5 22.98/0.7837/0.2143
+Proposed 1 26.90/0.8565/0.1316
+Improvement +3.92/+0.0728/-0.0827
+B.4 Detail Ablation
+We provide a module-wise detailed ablation study of the proposed DAIR in Table 15.
+This table illustrates the contributions of key components to the overall restoration performance.
+Each row represents a different configuration, with tick marks () indicating the inclusion of specific components and crosses () indicating their exclusion.
+The best results are highlighted in bold red.
+Notably, the full DAIR model demonstrates superior performance across all metrics, validating the effectiveness of the integrated components.
+Table 15: Ablation study on key components.
+Method LP LF DM 3WD Denoise Desnowing Derain Lowlight
+PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
+Base Model 23.42/0.724 20.76/0.770 22.70/0.770 13.99/0.511
+Base + LP 24.50/0.742 20.75/0.778 22.46/0.766 14.23/0.524
+Base + LP + LF 25.83/0.780 21.21/0.801 22.23/0.755 15.04/0.547
+Base + LP + LF + DM 26.80/0.816 21.97/0.844 22.79/0.761 15.52/0.554
+DAIR w/o Latent Prior 26.07/0.886 26.83/0.831 22.76/0.774 15.00/0.585
+DAIR 29.11/0.897 31.59/0.958 30.21/0.907 16.68/0.664
+Appendix C Additional Comparison with SOTA Methods
+C.1 Comparison on Three-task Settings
+Our primary motivation is to automate the prompting process and reframe AIR for practical, real-world applications.
+In real-world scenarios, degraded scenes are often highly complex and frequently involve overlapping degradations.
+To demonstrate the effectiveness of our method in addressing both homogeneous and heterogeneous degradations, we evaluate its practicality in a challenging six-task heterogeneous setting comprising complex datasets (LSD for real-world lowlight (sharif2025illuminatingdarknesslearningenhance), random denoising instead of fixed denoising, heavy rain streak, etc.).
+We refer to these tasks as a common task in the manuscript.
+Furthermore, we incorporate a compound 5D task, where degradations are similar and overlapping (e.g., Haze + Snow, Haze + Lowlight + Snow).
+Thus, we can illustrate the limitations of existing work and clearly articulate our motivation.
+However, in recent times, many methods, including PromptIR, ADAIR, DFPIR, Perceive-IR, etc., have leveraged three task settings to illustrate restoration performance.
+Therefore, to establish a clear positioning of our proposed method relative to existing models, we benchmarked it under similar settings.
+Table 16 compares our DAIR with existing methods.
+Our proposed method achieves notable improvements of +1.13 dB in dehazing and +0.28 dB on average over the best baselines.
+These results underscore the effectiveness of DAIR in enhancing restoration quality while being significantly lightweight.
+Table 16: Comparison of image restoration methods across different tasks and noise levels. Best results in bold red, second best underlined, and increment over best baseline highlighted in blue.
+Method Dehazing Deraining Denoising (CBSD68) Average
+SOTS Rain100L $\sigma=15$ $\sigma=25$ $\sigma=50$
+PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
+PromptIR 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913
+Restormer 30.43/0.975 36.55/0.975 33.84/0.931 31.18/0.885 27.90/0.790 31.98/0.911
+DFPIR 31.87/0.980 38.65/0.982 34.14/0.935 31.47/0.893 28.25/0.806 32.88/0.919
+ADAIR 31.06/0.980 38.64/0.983 34.12/0.935 31.45/0.892 28.19/0.802 32.69/0.918
+DAIR(Ours) 32.80/0.981 38.86/0.983 34.31/0.934 31.60/0.893 28.22/0.807 33.16/0.920
+Improvement +0.93/+0.001 +0.22/+0.000 +0.17/-0.001 +0.13/+0.000 -0.03/+0.001 +0.28/+0.001
+C.2 Common Task Details
+C.2.1 Low-light
+We utilize the LSD dataset (sharif2025illuminatingdarknesslearningenhance), collected in uncontrolled low-light settings, offering diverse indoor and outdoor scenes under varying conditions.
 Table 17 compares image restoration methods across extreme lowlight (under 50 Lux) and lowlight scenarios (50-200 lux), divided into indoor and outdoor subsets, using PSNR and SSIM metrics.
@@
 PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
-Uformer Wang et al. (2022) 11.98/0.5278 13.24/0.4628 10.39/0.5535 12.79/0.4145
-Restormer Zamir et al. (2022) 14.99/0.6343 16.35/0.5828 16.07/0.7121 13.90/0.5429
-AIRNet Li et al. (2022) 8.51/0.3678 9.80/0.3248 8.13/0.4585 9.05/0.3107
-PromptIR Vaishnav et al. (2023) 14.74/0.5902 16.10/0.5038 14.59/0.6404 13.48/0.4154
-DiffUIR Zheng et al. (2024) 9.77/0.5185 11.86/0.3713 9.77/0.5185 11.29/0.3193
-ADAIR Cui et al. (2024) 12.92/0.5456 14.46/0.4783 10.93/0.5576 12.72/0.4079
-DFPIR Tian et al. (2025a) 13.32/0.5643 15.68/0.4815 15.63/0.6599 14.74/0.3871
+Uformer 11.98/0.5278 13.24/0.4628 10.39/0.5535 12.79/0.4145
+Restormer 14.99/0.6343 16.35/0.5828 16.07/0.7121 13.90/0.5429
+AIRNet 8.51/0.3678 9.80/0.3248 8.13/0.4585 9.05/0.3107
+PromptIR 14.74/0.5902 16.10/0.5038 14.59/0.6404 13.48/0.4154
+DiffUIR 9.77/0.5185 11.86/0.3713 9.77/0.5185 11.29/0.3193
+ADAIR 12.92/0.5456 14.46/0.4783 10.93/0.5576 12.72/0.4079
+DFPIR 13.32/0.5643 15.68/0.4815 15.63/0.6599 14.74/0.3871
 DAIR (Ours) 17.63/0.7080 17.94/0.6486 17.13/0.7180 14.77/0.5898
 Improvement +2.64/+0.0737 +1.59/+0.0658 +1.06/+0.0059 +0.87/+0.0469
-B.1.2 Denoising
+C.2.2 Denoising
 Our experiment assesses the generalization of image denoising models under variable noise conditions, addressing limitations of prior works that train on fixed noise levels, leading to overfitting.
-We retrained baseline methods and our DAIR model using random Gaussian noise on DIV2K Agustsson & Timofte (2017) and evaluated on unseen datasets, BSD100 Martin et al.
-(2001) and Urban100 Kim et al.
-(2016).
+We retrained baseline methods and our DAIR model using random Gaussian noise on DIV2K (agustsson2017div2k) and evaluated on unseen datasets, BSD100 (martin2001bsd) and Urban100 (kim2016vdsr).
 Table 18 illustrate performance comparison across these datasets and noise levels ($\sigma$ = 15, 25, 50).
-Notably, baseline models showed inconsistent performance; Restormer struggled at higher noise, DiffUIR Zheng et al.
-(2024) excelled on BSD100 at $\sigma$=50 but failed elsewhere, on six-task common AIR settings.
-Notably, MP-based methods, such as PromptIR Vaishnav et al.
-(2023), underperformed in the absence of precise noise-level prompts.
+Notably, baseline models showed inconsistent performance; Restormer struggled at higher noise, DiffUIR (zheng2024selective) excelled on BSD100 at $\sigma$=50 but failed elsewhere, on six-task common AIR settings.
+Notably, MP-based methods, such as DFPIR (tian2025degradation), underperformed in the absence of precise noise-level prompts.
 This inconsistent performance, particularly among MP-based methods, underscores their inability to effectively segregate degradations, even on homogeneous tasks.
@@
 PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
-Uformer Wang et al. (2022) 31.71/0.9265 29.16/0.8753 25.46/0.7351 33.28/0.9219 30.61/0.8635 33.28/0.9219 31.52/0.9387 29.13/0.8944 25.15/0.7664
-Restormer Zamir et al. (2022) 27.38/0.8567 23.44/0.6825 16.62/0.3646 28.32/0.8385 23.73/0.6373 31.59/0.8937 28.11/0.8873 23.67/0.7304 16.52/0.4315
-AIRNet Li et al. (2022) 32.91/0.9377 30.52/0.9095 27.07/0.8307 34.20/0.9333 31.83/0.8986 27.03/0.7006 33.19/0.9506 31.00/0.9324 27.46/0.8665
-PromptIR Vaishnav et al. (2023) 28.84/0.8464 25.90/0.7306 21.45/0.5164 30.04/0.8295 27.03/0.7007 30.61/0.8636 29.23/0.8755 26.05/0.7734 21.23/0.5685
-DiffUIR Zheng et al. (2024) 19.92/0.8065 15.95/0.6774 14.34/0.5435 21.78/0.7909 17.41/0.6917 31.99/0.9122 18.28/0.7967 14.86/0.6701 13.67/0.5496
-ADAIR Cui et al. (2024) 32.51/0.9412 30.10/0.9075 26.80/0.8174 34.06/0.9354 31.59/0.8937 14.07/0.4531 32.74/0.9514 30.33/0.9233 26.72/0.8420
-DFPIR Tian et al. (2025a) 22.13/0.8095 21.57/0.7585 20.43/0.6234 21.77/0.8164 21.47/0.7535 21.48/0.7533 22.36/0.8182 21.80/0.7759 21.48/0.7533
+Uformer 31.71/0.9265 29.16/0.8753 25.46/0.7351 33.28/0.9219 30.61/0.8635 33.28/0.9219 31.52/0.9387 29.13/0.8944 25.15/0.7664
+Restormer 27.38/0.8567 23.44/0.6825 16.62/0.3646 28.32/0.8385 23.73/0.6373 31.59/0.8937 28.11/0.8873 23.67/0.7304 16.52/0.4315
+AIRNet 32.91/0.9377 30.52/0.9095 27.07/0.8307 34.20/0.9333 31.83/0.8986 27.03/0.7006 33.19/0.9506 31.00/0.9324 27.46/0.8665
+PromptIR 28.84/0.8464 25.90/0.7306 21.45/0.5164 30.04/0.8295 27.03/0.7007 30.61/0.8636 29.23/0.8755 26.05/0.7734 21.23/0.5685
+DiffUIR 19.92/0.8065 15.95/0.6774 14.34/0.5435 21.78/0.7909 17.41/0.6917 31.99/0.9122 18.28/0.7967 14.86/0.6701 13.67/0.5496
+ADAIR 32.51/0.9412 30.10/0.9075 26.80/0.8174 34.06/0.9354 31.59/0.8937 14.07/0.4531 32.74/0.9514 30.33/0.9233 26.72/0.8420
+DFPIR 22.13/0.8095 21.57/0.7585 20.43/0.6234 21.77/0.8164 21.47/0.7535 21.48/0.7533 22.36/0.8182 21.80/0.7759 21.48/0.7533
 DAIR (Ours) 33.08/0.9541 30.62/0.9252 27.25/0.8555 34.48/0.9468 31.99/0.9122 31.83/0.8988 33.23/0.9623 30.82/0.9395 27.30/0.8796
 Improvement +0.17/+0.0164 +0.10/+0.0157 +0.18/+0.0248 +0.28/+0.0135 +0.16/+0.0136 -0.16/-0.0134 +0.04/+0.0117 -0.18/+0.0071 -0.16/+0.0131
-B.2 More Visual Results
+C.3 More Visual Results
 Fig.
-7 presents visual results for denoising, deraining, and dehazing tasks, comparing our method with ADAIR Cui et al.
-(2024), DFPIR Tian et al.
-(2025a), degraded inputs, and reference images.
+9 presents visual results for denoising, deraining, and dehazing tasks, comparing our method with ADAIR (cui2024adair), DFPIR (tian2025degradation), degraded inputs, and reference images.
 These examples demonstrate that our approach consistently achieves clearer and more accurate restorations, closely matching the reference images and outperforming baselines across all degradation types.
@@
 Fig.
-8 presents more visual results for compound degradations (e.g., low-light + haze + rain, haze + rain), comparing our method with ADAIR Cui et al.
-(2024), DFPIR Tian et al.
-(2025a), degraded inputs, and reference images.
+10 presents more visual results for compound degradations (e.g., low-light + haze + rain, haze + rain), comparing our method with ADAIR (cui2024adair), DFPIR (tian2025degradation), degraded inputs, and reference images.
 Our approach consistently delivers clearer and more accurate restorations, outperforming baselines in challenging compound degradation settings.
@@
 Our method consistently delivers clearer and more accurate outputs, closely matching the reference images and outperforming ADAIR and DFPIR.
-B.3 Inference Benchmarking
+C.4 Inference Benchmarking
 Table 19: Inference speed comparison across different image restoration models. Best results in bold red, second best underlined.
@@
 Time (ms)↓ FPS↑ Throughput (MP/s)↑ Time (ms)↓ FPS↑ Throughput (MP/s)↑ Time (ms)↓ FPS↑ Throughput (MP/s)↑
-Uformer Wang et al. (2022) 31.07 32.19 0.53 58.08 17.22 1.13 203.83 4.91 1.29
-PromptIR Vaishnav et al. (2023) 65.62 15.24 0.25 180.31 5.55 0.36 707.17 1.41 0.37
-ADAIR Cui et al. (2024) 71.01 14.08 0.23 191.50 5.22 0.34 732.94 1.36 0.36
-DFPIR (with CLIP) Tian et al. (2025a) 70.08 14.27 0.23 184.33 5.43 0.36 714.20 1.40 0.37
-DFPIR (dummy text) Tian et al. (2025a) 67.88 14.73 0.24 184.12 5.43 0.36 714.38 1.40 0.37
+Uformer 31.07 32.19 0.53 58.08 17.22 1.13 203.83 4.91 1.29
+PromptIR 65.62 15.24 0.25 180.31 5.55 0.36 707.17 1.41 0.37
+ADAIR 71.01 14.08 0.23 191.50 5.22 0.34 732.94 1.36 0.36
+DFPIR (with CLIP) 70.08 14.27 0.23 184.33 5.43 0.36 714.20 1.40 0.37
+DFPIR (dummy text) 67.88 14.73 0.24 184.12 5.43 0.36 714.38 1.40 0.37
 DAIR (Ours) 20.55 48.67 0.80 55.20 18.12 1.19 220.92 4.53 1.19
@@
 Our DAIR model consistently achieves the fastest processing times, highest FPS, and superior throughput across various image sizes, outperforming all baseline AIR models.
-Notably, DAIR matches or surpasses the single-task model like Uformer (43.86G) Wang et al.
-(2022), which is significantly lighter compared to the existing AIR methods.
+Notably, DAIR matches or surpasses the single-task model like Uformer (43.86G) (wang2022uformer), which is significantly lighter compared to the existing AIR methods.
 These results highlight DAIR’s efficiency, demonstrating its ability to deliver SOTA restoration quality while maintaining faster speed and practicality for real-world applications, even on modest hardware setups.
-B.4 Downstream Vision Tasks (Object Detection)
-A primary objective of AIR is enhancing downstream vision task performance Cui et al.
-(2024); Jiang et al.
-(2025), yet existing AIR methods lack benchmarking on such tasks due to insufficient annotated datasets.
-To address this critical gap, we created a comprehensive evaluation framework by annotating 3,000 test images across six common degradations (lowlight Sharif et al.
-(2025), deraining Fu et al.
-(2017), dehazing Li et al.
-(2018), desnowing Li et al.
-(2020), denoising Agustsson & Timofte (2017), and deblurring Nah et al.
-(2017)).
+Appendix D More Results from Real-world
+D.1 Downstream Vision Tasks (OD)
+A primary objective of AIR is enhancing downstream vision task performance (cui2024adair; jiang2025survey), yet existing AIR methods lack benchmarking on such tasks due to insufficient annotated datasets.
+To address this critical gap, we created a comprehensive evaluation framework by annotating 3,000 test images across six common degradations (lowlight (sharif2025illuminatingdarknesslearningenhance), deraining (fu2017deep), dehazing (li2018benchmarking), desnowing (li2020desnownet), denoising (agustsson2017div2k), and deblurring (nah2017deep)).
 Our annotation process leveraged the paired nature of these datasets by first annotating ground truth reference images and transferring these annotations to their corresponding corrupted counterparts, ensuring consistency and reliability.
-We then enhanced these degraded images using several baseline AIR restoration methods and systematically evaluated their performance using the YOLOv12-Large Tian et al.
-(2025b) object detection model.
-This novel benchmarking approach provides the first quantitative assessment of how different restoration techniques impact object detection performance, offering valuable insights for developing AIR methods optimized for real-world computer vision applications.
+We then enhanced these degraded images using several baseline AIR restoration methods and systematically evaluated their performance using the YOLOv12-Large (tian2025yolov12) OD model.
+This novel benchmarking approach provides the first quantitative assessment of how different restoration techniques impact OD performance, offering valuable insights for developing AIR methods optimized for real-world computer vision applications.
 Table 20: Performance comparison of object detection models. Best results in bold red, second best underlined.
@@
 IoU=0.50:0.95 IoU=0.50 IoU=0.75 Small Medium Large
-Uformer Wang et al. (2022) 28.0 30.2 29.4 36.5 26.9 27.3
-Restformer Zamir et al. (2022) 31.7 34.1 33.2 37.5 28.7 32.4
-AIRNet Li et al. (2022) 30.2 32.7 31.8 35.8 26.4 30.3
-PromptIR Vaishnav et al. (2023) 30.9 33.3 31.9 41.8 28.6 30.6
-ADAIR Cui et al. (2024) 28.9 31.2 30.2 38.4 28.8 28.1
-DFPIR Tian et al. (2025a) 27.9 30.0 29.3 26.6 25.5 29.8
+Uformer 28.0 30.2 29.4 36.5 26.9 27.3
+Restormer 31.7 34.1 33.2 37.5 28.7 32.4
+AIRNet 30.2 32.7 31.8 35.8 26.4 30.3
+PromptIR 30.9 33.3 31.9 41.8 28.6 30.6
+ADAIR 28.9 31.2 30.2 38.4 28.8 28.1
+DFPIR 27.9 30.0 29.3 26.6 25.5 29.8
 DAIR (Our) 34.9 37.3 36.4 44.9 30.8 35.1
 Improvement +3.2 +3.2 +3.2 +3.1 +2.0 +2.7
-Table 10 presents a comprehensive performance comparison of various image restoration methods for object detection tasks.
-Our proposed DAIR method demonstrates superior performance across all evaluation metrics, achieving the highest Average Precision (AP) scores of 34.9% at IoU=0.50:0.95, 37.3% at IoU=0.50, and 36.4% at IoU=0.75, surpassing the second-best method (Restformer Zamir et al.
-(2022)) by a significant margin of 3.2 percentage points.DAIR also excels in detecting objects of varying sizes, achieving notably high AP scores on small (44.9%, outperforming PromptIR’s 41.8%), medium (30.8%), and large (35.1%) objects.
-The consistent improvements across all metrics highlight DAIR’s effectiveness in enhancing image quality for downstream object detection, with an average improvement of 3.2 percentage points over SOTA methods.
-These results demonstrate DAIR’s robust capability to address diverse degradation scenarios while preserving essential visual information for accurate object detection.
-Appendix C Unseen Task Generalization
+Table 20 presents a comprehensive performance comparison of various image restoration methods for OD tasks.
+Our proposed DAIR method demonstrates superior performance across all evaluation metrics, achieving the highest Average Precision (AP) scores of 34.9% at IoU=0.50:0.95, 37.3% at IoU=0.50, and 36.4% at IoU=0.75, surpassing the second-best method (Restformer (zamir2022restormer)) by a significant margin of 3.2 percentage points.DAIR also excels in detecting objects of varying sizes, achieving notably high AP scores on small (44.9%, outperforming PromptIR’s 41.8%), medium (30.8%), and large (35.1%) objects.
+The consistent improvements across all metrics highlight DAIR’s effectiveness in enhancing image quality for downstream OD, with an average improvement of 3.2 percentage points over SOTA methods.
+These results demonstrate DAIR’s robust capability to address diverse degradation scenarios while preserving essential visual information for accurate OD.
+D.2 Unseen Task Generalization
 A key motivation behind the proposed latent-prior DAIR framework is its ability to generalize to diverse, previously unseen degradations.
-To validate this capability, we extensively evaluate DAIR under cross-domain and out-of-distribution scenarios, including underwater image enhancement Peng & Cosman (2019), medical image perceptual enhancement Uhlen et al.
-(2010), medical image denoising Rezvantalab et al.
-(2018), real-world denoising Abdelhamed et al.
-(2018), and unseen real-world low-light enhancement Sharif et al.
-(2025).
-Notably, all evaluations are performed using a DAIR model trained only on common degradations, without any TS fine-tuning or external guidance.
-In addition to the visual results, we summarized the quantitative score with no-reference metrics: NIQE Mittal et al.
-(2013), MUSIQ Ke et al.
-(2021), and LPIPS Zhang et al.
-(2018b).
-C.1 Real-World Denoising
-Table.
-8 and Figure 9 illustrate the performance of the proposed method on real-world noisy images Abdelhamed et al.
-(2018).
+To validate this capability, we extensively evaluate DAIR under cross-domain and out-of-distribution scenarios, including underwater image enhancement (peng2017uieb), medical image perceptual enhancement (uhlen2010towards), medical image denoising (rezvantalab2018dermatologist), real-world denoising (abdelhamed2018ssid), and unseen real-world low-light enhancement (sharif2025illuminatingdarknesslearningenhance).
+D.2.1 Comparison with existing Methods.
+We also compared existing methods against DAIR on unseen real-world degradation.
+We benchmarked all models trained on common degradation without any task-specific fine-tuning.
+The quantitative scores are summarized using no-reference metrics: NIQE mittal2013niqe, MUSIQ ke2021musiq, and LPIPS zhang2018lpips.
+Table 21 presents a performance comparison on no-reference quality metrics.
+Our proposed method, DAIR, achieves the best results in most categories, demonstrating significant improvements, including a reduction of -0.78 in NIQE for unseen Real-Lowlight and an increase of +5.08 in MUSIQ.
+Furthermore, DAIR excels in Real-Noise and Real Perceptual Enhancement tasks, highlighting its effectiveness in enhancing image quality across a range of applications.
+Table 21: Performance comparison across different image restoration tasks using no-reference quality metrics.
+Best results in bold red, second best underlined, and increment over best performing method highlighted in blue.
+Method Real-Lowlight Real-Noise Real Underwater Real Perceptual Enhancement Medical Denoising Average
+NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓
+Input 5.33/47.30/27.17 4.44/62.56/5.55 3.83/52.35/8.80 13.94/37.06/56.26 17.06/21.81/65.10 8.92/44.42/32.58
+PromptIR 5.31/42.58/39.07 5.77/61.11/27.38 3.98/54.00/17.45 7.59/42.62/38.14 25.60/18.72/78.09 9.65/43.81/40.03
+DFPIR 5.50/45.81/21.83 4.88/59.72/16.59 4.06/52.16/8.97 9.55/43.13/15.52 16.04/24.27/64.71 8.01/45.02/25.52
+ADAIR 5.53/46.71/35.77 4.38/63.69/16.30 3.63/55.50/13.29 11.08/40.27/33.54 9.65/24.34/43.36 6.85/46.10/28.45
+DAIR (Ours) 4.75/51.79/21.97 4.24/69.37/16.94 3.61/55.32/15.50 5.42/45.94/34.91 8.22/28.53/24.55 5.25/50.19/25.97
+Improvement -0.78/+5.08/+0.14 -0.14/+5.68/+0.35 -0.02/-0.18/+6.70 -2.17/+2.81/+19.39 -1.43/+4.19/-18.81 -1.60/+4.09/+0.45
+D.2.2 Visual results
+Figure 11 illustrate the performance of the proposed method on real-world noisy images (abdelhamed2018ssid).
 In well-exposed scenes, DAIR performs only denoising, preserving natural brightness and contrast.
-However, when inputs exhibit low-light characteristics (as commonly observed in SSID-like datasets Abdelhamed et al.
-(2018)) along with noise, the latent descriptor guides DAIR to jointly suppress noise and enhance brightness.
+However, when inputs exhibit low-light characteristics (as commonly observed in SSID-like datasets (abdelhamed2018ssid)) along with noise, the latent descriptor guides DAIR to jointly suppress noise and enhance brightness.
 Thus, DAIR adaptively performs denoising and enhancement without prompts or handcrafted rules.
-Table 11: Image Quality Metrics for Real-world Noise Dataset. Best results in bold red.
-Method Image Quality Metrics
-NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
-Input 5.57 61.31 0.81
-Enhanced (Our) 4.24 69.37 0.80
-Improvement +1.33 +8.06 +0.01
-Figure 9: Real-world denoising results.
-Left: low-light noisy scenes where DAIR jointly performs denoising and LLIE automatically, without user input.
-Right: well-lit, noisy scenes where DAIR performs only denoising.
-C.2 Underwater Image Enhancement
-We further evaluate DAIR on underwater images, a domain unseen during training.
-As shown in Figure 10, the model adapts to varying underwater degradations, including haze-like scattering (left examples) and low exposure (right examples).
-Despite never being trained on underwater data, DAIR enhances visibility and color fidelity, demonstrating strong cross-domain generalization driven by its latent prior representation.
-Table 12 further details DAIR’s performance on underwater image enhancement.
-Figure 10: Underwater image enhancement. DAIR generalizes to haze-like and low-exposed underwater degradations, improving contrast and color balance without domain-specific training.
-Table 12: Image Quality Metrics for Underwater Image Enhancement. Best results in bold red.
-Method Image Quality Metrics
-NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
-Input 3.83 53.35 0.70
-Enhanced (Our) 3.78 54.05 0.70
-Improvement +0.05 +0.70 +0.00
-C.3 Unseen Low-Light Image Enhancement
-We evaluate DAIR on real-world low-light images captured with unknown hardware and distributions (e.g., LSD-U Sharif et al.
-(2025)) unseen during training.
-Figure 11 and Table 13 demonstrate that the latent prior effectively encodes degradation properties, enabling robust out-of-distribution LLIE.
-Figure 11: Unseen real-world low-light enhancement. DAIR can enhance illumination and structure in previously unseen low-light scenes without artifacts or over-enhancing the scenes.
-Table 13: Image Quality Metrics for Real-world Lowlight Enhancement. Best results in bold red.
-Method Image Quality Metrics
-NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
-Input 4.47 62.33 0.81
-Enhanced (Our) 4.14 62.51 0.81
-Improvement +0.33 +0.18 +0.00
-C.4 Medical Image Perceptual Enhancement
-We also study cross-domain perceptual enhancement on microscopic medical images Uhlen et al.
-(2010), which often suffer from poor contrast and weak structural saliency Sharif et al.
-(2022).
-Despite having no exposure to this modality during training, DAIR enhances perceptual quality by improving local contrast and clarity while preserving natural textures (Figure 12).
-These results demonstrate that the latent prior generalizes to structurally distinct domains.
-Table 14 presents the corresponding quantitative improvements.
-Figure 12: Perceptual enhancement of microscopic medical images. DAIR improves contrast and structural visibility in unseen medical imagery without TS training.
-Table 14: Image Quality Metrics for Perceptual Enhancement. Best results in bold red.
-Method Image Quality Metrics
-NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
-Input 13.94 37.06 0.63
-Enhanced (Our) 9.68 41.50 0.79
-Improvement -4.26 +4.44 -0.16
-C.5 Medical Image Denoising
-Finally, we evaluate DAIR on medical image denoising using a dataset not seen during training, specifically dermatoscopic imagery Rezvantalab et al.
-(2018).
-As shown in Figure 13, DAIR effectively suppresses noise while preserving diagnostically relevant textures and edges, without requiring manual prompting.
-Table 15 illustrates the performance of DAIR on unseen medical image denoising.
-Figure 13: Medical image denoising. DAIR reduces noise and preserves fine structural details in previously unseen medical data, operating without guidance or fine-tuning.
-Table 15: Image Quality Metrics for Medical Noise Reduction. Best results in bold red.
-Method Image Quality Metrics
-NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
-Input 16.99 21.93 0.82
-Enhanced (Our) 12.64 25.17 0.70
-Improvement +4.35 +3.24 +0.12
-Appendix D Limitation and Future Scope
-Despite the significant improvement of our proposed DAIR method over existing approaches on both common and compound degradations, as well as its performance on unseen tasks, we observed certain limitations.
-Figure 14 illustrates some failure cases of our method on unseen tasks.
+Figure 11: Real-world unseen degradation restoration with DAIR. Our method can restore unseen out-of-distribution datasets without needing any fine tuning.
+Similarly, our method can significantly improve the visual quality of unseen diverse real-world degradation, as shown in Fig. 12
+Figure 12: Real-world unseen degradation restoration with DAIR. Our method can restore unseen out-of-distribution datasets without needing any fine tuning.
+D.3 Real-world Compound Degradation
+We evaluated our method on real-world compound degradation (i.e., LOL-Blur (zhou2022lednet).
+Table 22 presents a quantitative comparison of image restoration models on synthesized and real datasets.
+The results demonstrate that our proposed method, DAIR, outperforms existing models in all evaluated metrics.
+Notably, DAIR achieves the highest PSNR and SSIM values, as well as the lowest LPIPS and NIQE scores, indicating superior restoration quality compared to ADAIR.
+DAIR achives a PSNR of 24.10 and an SSIM of 0.8532 few-shot fine tunning.
+Fig.
+13 further demonstrates the practicality of our method through a visual comparison.
+Table 22: Quantitative comparison on synthesized and real datasets. Best results in bold red.
+Model Training Synthesized Real
+PSNR↑ SSIM↑ LPIPS↓ NIQE↓ MUSIQ↑ BRISQUE↓
+Low (input) - 8.74 0.4255 0.4943 7.04 32.18 36.22
+ADAIR Zero-shot 11.77 0.5528 0.4057 6.80 33.62 32.57
+DAIR Zero-shot 13.35 0.6221 0.3393 5.14 46.39 28.58
+ADAIR Fine-tune 22.83 0.7811 0.2406 6.80 33.62 32.57
+DAIR Fine-tune 24.10 0.8532 0.1503 4.70 49.58 27.74
+Figure 13: Real-world compound degradation restoration with DAIR.
+Appendix E Limitation and Future Scope
+Despite the significant improvements of our DAIR method over existing approaches for both common and compound degradations, we identified some limitations.
+In real-world compound scenes, such as those in the LOL-Blur dataset(zhou2022lednet), our method can produce artifacts.
+Figure 14 shows failure cases on unseen tasks.
 Specifically, in extreme low-light conditions (under 5 lux), artifacts were observed in homogeneous spatial regions, such as cloud-free skies.